### v1

# A Simple DropConnect Approach to Transfer-based Targeted Attack

Abstract
We study the problem of transfer-based black-box attack, where adversarial samples generated using a single surrogate model are directly applied to target models. Compared with untargeted attacks, existing methods still have lower Attack Success Rates (ASRs) in the targeted setting, i.e., the obtained adversarial examples often overfit the surrogate model but fail to mislead other models. In this paper, we hypothesize that the pixels or features in these adversarial examples collaborate in a highly dependent manner to maximize the success of an adversarial attack on the surrogate model, which we refer to as perturbation co-adaptation. Then, we propose to Mitigate perturbation Co-adaptation by DropConnect (MCD) to enhance transferability, by creating diverse variants of surrogate model at each optimization iteration. We conduct extensive experiments across various CNN- and Transformer-based models to demonstrate the effectiveness of MCD. In the challenging scenario of transferring from a CNN-based model to Transformer-based models, MCD achieves 13% higher average ASRs compared with state-of-the-art baselines. MCD  boosts the performance of self-ensemble methods by bringing in more diversification across the variants while reserving sufficient semantic information for each variant. In addition, MCD attains the highest performance gain when scaling the compute of crafting adversarial examples.
footnotetext: Preprint. Work in progress.
1 Introduction
In recent years, deep neural networks have become prevalent in computer vision applications [15, 17, 9], but are highly vulnerable to maliciously crafted inputs, which are called adversarial examples [34, 11]. A major concern is their transferability, that is, adversarial examples generated using a white-box model can be directly used to fool other black-box models, without any access to their architectures, parameters or gradients [13]. Since this type of attacks, usually referred to as transfer-based attacks, do not require any interaction with the target model, they pose severe security risks to real-world machine learning systems. Therefore, studying effective transfer-based attack methods is crucial to understanding the vulnerabilities, and in turn enhancing the robustness of machine learning models.
This paper focuses on targeted transfer-based attacks with a single surrogate model, where the goal is to deceive black-box models to classify input images into a specific incorrect category. Due to the high complexity of decision boundaries, existing methods still have noticeably lower Attack Success Rates (ASRs) in the targeted setting than in the untargeted [47, 2]. A key observation is that the generated adversarial examples tend to overfit the surrogate model but fail to generalize to other models. To improve transferability, various strategies have since been proposed. Noticing that multiple surrogate models can help enhance transferability but in practice finding proper models for the same task is not easy [7, 26, 25, 46], model self-ensemble [18, 45, 22, 28] tries to create multiple models from only one accessible model, which is in accordance with our setting. Input transformation [44, 48, 1, 42, 40, 2] applies different transformations to inputs and diversify input patterns in order to reduce overfitting. A notable method is Clean Feature Mixup (CFM) [2], which randomly mixes high-level features with shuffled clean features and has achieved state-of-the-art performance. Despite these progresses, there is still much room for further improvement.
In this work, we hypothesize that there is a key aspect underlying the present overfitting issue: the pixels in many adversarial examples, obtained from existing transferred-based methods, collaborate in a highly dependent manner to maximize the success of adversarial attacks on the surrogate model, which we refer to as perturbation co-adaptation. Similar to the co-adaption issue of units in training neural networks [33], we believe that perturbation co-adaptation would lead to overly complex patterns that can fool the surrogate model but fail on other target models. How to further alleviate this co-adaptation issue would be a key to improving the transferability of adversarial examples over existing transfer-based attack methods.
To this end, we propose to Mitigate perturbation Co-adaptation by DropConnect (MCD), motivated by the classic Dropout technique that can effectively prevent units from co-adaptation during training neural networks [39]. Rather than directly applying Dropout to the input or the perturbation, our method extends random masking to 1) high-level features that contain richer semantic information, and 2) model parameter space for more flexible and diverse feature disturbances for perturbation. Fig. 1 illustrates the proposed method. At each optimization iteration, MCD  creates diverse variants of the surrogate model with masked parameters, which are then used to update the adversarial examples. In this sense, the final algorithm can be regarded as a self-ensemble method. Compared with other related self-ensemble methods [45, 22], MCD brings in more diversification across the variants while reserving sufficient semantic information for each variant, hence improves the transfer attack performance. In addition, MCD can be easily implemented into all layers in the surrogate model, with little comprise in terms of attack performance. A more concrete comparison is left to Sec. 4.3.
Figure 1: An illustration of the proposed method MCD. We apply MCD to selected layers in the surrogate model to create multiple and diverse variants at each optimization iteration.
We evaluate the proposed method across various CNN- and Transformer-based target models, as well as against several defense methods. The experimental results show that MCD outperforms other state-of-the-art methods. In particular, in the challenging scenario of transferring from CNN-based model to Transformer-based models, MCD achieves 13.0% and 17.9% higher average ASRs with ResNet-50 [15] and DenseNet-121 [17] as surrogate models, respectively. Moreover, MCD achieves the highest performance gain when scaling the compute for crafting adversarial examples. Specifically, with ResNet-50 as surrogate model, increasing the optimization iterations from 300 to 500 and number of forward-backward passes per iteration from 1 to 5 boosts the average ASR by 15.9%.
To summarize, our main contributions are as follows:
  • We show the perturbation co-adaptation issue in existing transfer-based attack methods in the targeted setting. Reducing this co-adaptation issue would in turn enhance transferability of adversarial examples.
  • We propose MCD to improve the transferability, by creating diverse variants of surrogate model during optimization. MCD brings in more diversification across the variants while reserving sufficient semantic information for each variant, and is also simple to implement.
  • Extensive experiments across various surrogate and target models validate the superior performance of MCD over other methods. Moreover, MCD benefits the most when we scale the compute of crafting adversarial examples.
2 Related Works
In this section, we provide a brief review of related works focusing on transfer-based attack methods.
One of the most fundamental attack methods is Fast Gradient Sign Method (FGSM) [11], which uses the direction of gradient to craft adversarial examples. Iterative-FGSM (I-FGSM) [20] extends FGSM into an iterative framework to enhance the attack performance. However, while the obtained adversarial examples achieve high success rates for white-box attacks, they do not perform well when transferred to other black-box models. To improve transferability, many methods have been proposed, mainly from the following perspectives: input transformation, gradient stabilization, loss function refinement, and model ensemble.
The first class applies transformations to the input images, thereby diversifying the input patterns [44, 48, 8, 1, 42, 40, 2]. Diverse Inputs (DI) [44] is one of the representative methods. At each iteration, it applies random and differentiable transformations (e.g., random resizing) to the input image and then maximizes the loss function w.r.t. the transformed inputs. Resized Diverse Inputs (RDI) [48] extends DI by resizing the transformed image back to its original size. The Translation-Invariant (TI) attack method [8] optimizes perturbations across a set of translated images, making the adversarial example less sensitive to the surrogate model. Block Shuffle and Rotate (BSR) [40] divides the input image into blocks, followed by shuffling or rotation, and Object-based Diversity Inputs (ODI) [1] generates an adversarial example on a 3D object and leads the rendered image to being classified as the target class. Admix [42] mixes the input image with random samples from other classes. Clean Feature Mixup (CFM) [2] extends Admix to high-level feature space and introduces two types of competing noises to guide adversarial perturbations, achieving state-of-the-art performance in transfer-based targeted attacks.
The second class of methods stabilize the gradient updates to enhance transferability during adversarial example generation. Momentum Iterative FGSM (MI-FGSM) [7] incorporates a momentum term into I-FGSM, helping avoid local optima. The Variance-Tuned (VT) method [41] further improves stability by not only accumulating the gradients at current iteration but also incorporating the variance of gradients from previous iterations to adjust the current update. Scale-Invariant (SI) optimization [24] leverages the scale-invariance property of deep learning models by applying perturbations across multiple scaled copies of the input image, in order to reduce overfitting to the white-box model.
Another direction is to devise tailored loss functions, particularly for targeted attacks. The Po+Trip method [21] introduces Poincaré distance as a similarity measure and dynamically adjusts gradient magnitudes to mitigate the “noise curing” issue. The Logit method [47], which directly maximizes the logit output of the target class, alleviates the gradient vanishing problem and has demonstrated significant improvement w.r.t. transferability of targeted attack.
Additionally, surrogate model ensemble is also useful to enhancing transferability, where one utilizes the average of losses, predictions, or logits of multiple models to craft adversarial examples [7, 26]. A recent work [3] proposes the common weakness attack method composed of sharpness aware minimization and cosine similarity encourager, beyond the averaging. As shown in [18, 25], the transferability generally improves with an increasing number of surrogate classifiers. However, the number of surrogate models in practice is usually small and picking proper surrogate models for the same task is also not easy.
Closely related to the present work is self-ensemble [18, 45, 22, 28], which creates multiple models from only one surrogate model. Transfer-based self-ensemble (T-sea) [18] targets the task of object detection, and the self-ensemble method in [28] specifically considers vision Transformer as surrogate model and is denoted as SE-ViT in this paper. Ghost Network [22] perturbs surrogate model to create a set of new models and then samples one model from the set at each iteration. Masking Unimportant Parameters (MUP) [45] drops out unimportant parameters according to a predefined Taylor expansion-based metric.
3 Preliminary
Consider a classification task where the model is defined as a function $f:\mathbb{R}^{n}\to\mathcal{Y}$ that maps an input $x\in\mathbb{R}^{n}$ to a label in the set $\mathcal{Y}$ consisting of all the labels. Given a clean image $x$ with its true label $y\in\mathcal{Y}$, untargeted attacks aim to find an adversarial example $x_{\mathrm{adv}}\in\mathbb{R}^{n}$ that is similar to $x$ but misleads the model to produce an incorrect prediction, i.e. $f(x_{\mathrm{adv}})\neq y$. Here ‘similarity’ is usually measured by an $\ell_{p}$-norm, e.g., $\|x_{\mathrm{adv}}-x\|_{p}\leq\epsilon$ where $\epsilon>0$ is a predefined perturbation budget. For targeted attacks, the goal is to modify the model prediction to a particular target label $y_{\mathrm{tar}}$, that is, $f(x_{\mathrm{adv}})=y_{\mathrm{tar}}\neq y$. In this work, we will focus on targeted attacks.
Under the white-box setting where the model is fully accessible, adversarial example can be obtained through solving the following problem:
$\operatorname*{arg\,max}_{x_{\mathrm{adv}}}~{}\mathcal{L}\left(f(x_{\mathrm{adv}})\right),~{}~{}\text{s.t.}~{}~{}\|x_{\mathrm{adv}}-x\|_{p}\leq\epsilon,$ (1)
where $\mathcal{L}(\cdot)$ represents a loss function like cross-entropy loss. A strong and popular attack method to the above optimization problem is the Iterative Fast Gradient Sign Method (I-FGSM) [20] that leverages the sign of the gradient information. In particular, at each iteration $t$, the adversarial example is updated by adding a small perturbation:
$x_{\mathrm{adv}}^{t}=x_{\mathrm{adv}}^{t-1}+\alpha\cdot\mathrm{sign}\left(\nabla_{x_{\mathrm{adv}}^{t-1}}\mathcal{L}\left(f(x_{\mathrm{adv}}^{t-1})\right)\right),$ (2)
where $x_{\mathrm{adv}}^{0}=x$ and $\alpha>0$ is a step size. To make the generated adversarial example satisfy the perturbation budget constraint, a straightforward way is to project $x_{\mathrm{adv}}^{t}$ into the $\epsilon$-ball of $x$.
In the black-box settings, the gradient information is not available and we only have limited access to the target model. A large class of attack methods in this setting is transfer-based, which assumes that the adversarial examples exhibit transferability across different neural models; that is, adversarial examples generated on one model to mislead not only that model but also other models [34]. Thus, one can use the above updates (2) on an available white-box model, usually called surrogate model, to generate adversarial examples that are applied to the target model. To boost the performance of transfer-based methods, an effective approach is to perform some transformation $\mathcal{T}$ (like resizing [44, 48] and scaling [24]) to the input and then use $\nabla_{x_{\mathrm{adv}}^{t-1}}\mathcal{L}\left(f(\mathcal{T}(x_{\mathrm{adv}}^{t-1}))\right)$ as the gradient information.
However, the inherent differences w.r.t. model architecture, parameter setting, and training procedure between the surrogate and target models still limit the effectiveness of transfer-based methods on certain models and datasets. In the next section, we show that there is a key aspect that renders the current over-fitting issue of transfer-based methods.
4 Methodology
In this section, we first conduct a pilot study to show a key aspect of the overfitting issue in existing transfer-based methods and then propose a random masking based approach. Comparison with related methods is also discussed.
4.1 Motivation
We hypothesize that the pixels or features in the adversarial example, obtained through solving Problem (1), collaborate in a highly dependent manner to maximize the success of an adversarial attack on the surrogate model, which we refer to as perturbation co-adaptation. Even with state-of-the-art transfer-based attack methods, this issue is still a key aspect that can make the perturbation fail to mislead other models.
We now conduct a pilot study to verify our hypothesis. Since adversarial example is a composition of the given clean image and some perturbation, Problem (1) is equivalent to finding a perturbation that maximizes the loss function. In this sense, we perform random masking with different Dropout ratios to the obtained perturbation $\delta:=x_{\mathrm{adv}}-x$, to disrupt potential co-adaptations among pixels. Note that the masked version of perturbation always satisfies the $\ell_{p}$-norm constraint if the adversarial example already meets this constraint. We consider a targeted attack setting on an ImageNet-compatible dataset (see more details in Sec. 5.1), where adversarial examples are crafted using ResNet-50. The used attack methods include DI, ODI and CFM [44, 1, 2].
The results are shown in Fig. 2 (additional results can be found in Sec. B.1). We can observe an obvious decrease in ASRs as the masking ratio exceeds 20%, in the DI, ODI and CFM methods, indicating that the pixels in the obtained perturbation are likely to be heavily dependent to attain a successful targeted attack. Similar to the co-adaption issue in training neural networks [33], we believe that perturbation co-adaptation would lead to overly complex patterns that mislead surrogate model but fail on other models. As such, how to further alleviate this co-adaptation would be a key to improving the transferability of adversarial examples over existing transfer-based attack methods.
4.2 Mitigating Perturbation Co-adaptation by DropConnect
As per the pilot study, a direct approach to improving transferability would be to apply random masking to pixels in perturbation during the iterative updates. As shown in Sec. B.2, by encouraging perturbations to explore more diverse feature disturbances, this way indeed yields higher ASRs than the baseline DI and RDI methods in the targeted attack setting. However, it does not consistently achieve the best attack performance; a potential reason is that pixel-level masking may be less effective of capturing semantic information for image classification. Based on these initial results, it is tempting to extend this direct idea to craft more transferable adversarial examples.
[…]
(b) ResNet-50 $\to$ MobileNet-v2
Figure 2: ASRs(%) with increasing mask ratio.
The rest of this work is devoted to further mitigating the perturbation co-adaptation, by extending random masking to higher-level representations. The overall approach is to create multiple diverse variants of the surrogate model during optimization. We first consider the following questions:
  • Q1: how to achieve more flexible and diverse feature disturbances for perturbation than vanilla Dropout?
  • Q2: which intermediate representations or layers to perform the random masking?
For Q1, we observe that in typical machine learning problems where the goal is to optimize model parameters w.r.t. given training data, standard Dropout [33] is applied to the units or certain features. Finding adversarial example, as formulated in Problem (1), can be treated as optimizing perturbed image $x_{\mathrm{adv}}$ with fixed model parameters as ‘training data’ or ‘input features’. As analogy to Dropout, we can randomly drop out the model parameters connecting the units. This is indeed the DropConnect technique [39], a generalized version of Dropout by providing a more fine-grained way of regularizing the network during training. As such, we will focus on DropConnect in this paper.
For Q2, we notice that DropConnect is mainly effective in terms of linear layers [39]. We thus apply DropConnect to the weight and bias parameters of linear layers as well as the transformation parameters of normalization layers in the surrogate model. Both types of layers are widely used in mainstream architectures including Transformer [38]. Empirically, our ablation study in Sec. 5 validates the effectiveness of this choice, compared with convolutional layers.
We now present our attack method, Mitigating perturbation Co-adaptation by DropConnect (MCD), as summarized in Algorithm 1. Take linear layer for example. We perform independent random masking onto the weight and bias (when present) parameters using Bernoulli sampling. Specifically, for surrogate model $f$, let $W\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ denote the weight matrix and $b\in\mathbb{R}^{d_{\text{out}}}$ the bias vector associated with a selected linear layer. Here $d_{\text{in}}$ and $d_{\text{out}}$ represent the input and output dimensions, respectively. The corresponding masks for the weight matrix and bias are
$\mathcal{M}_{w}\sim\text{Bernoulli}(p_{w}),\quad\mathcal{M}_{b}\sim\text{Bernoulli}(p_{b}),$ (3)
where $\mathcal{M}_{w}\in\{0,1\}^{d_{\text{in}}\times d_{\text{out}}}$ and $\mathcal{M}_{b}\in\{0,1\}^{d_{\text{out}}}$ are the random masks, and $p_{w},p_{b}\in[0,1]$ are DropConnect probabilities. Then the masked parameters are computed as
$W_{\mathcal{M}}=\mathcal{M}_{w}\odot W,\quad b_{\mathcal{M}}=\mathcal{M}_{b}\odot b,$ (4)
where $\odot$ denotes the element-wise multiplication. For normalization layer, the same operation is applied similarly to the transformation parameters. The random masks are sampled for each selected layer.
Algorithm 1 Mitigating perturbation Co-adaptation by DropConnect (MCD)
1:Classifier $f$; clean image $x$; loss function $\mathcal{L}(\cdot)$; max iterations $T$; $\ell_{p}$ bound $\epsilon$; number of inferences per iteration $S$; DropConnect probabilities $p_{w}$, $p_{b}$; linear and normalization layers $\mathbb{L}$; input transformation $\mathcal{T}$.
2:Adversarial example $x_{\mathrm{adv}}$
3:$x_{\mathrm{adv}}^{0}\leftarrow x$
4:for $t=1\to T$ do
5:     for $s=1\to S$ do
6:         Obtain modified model $f_{\mathcal{M}}$ by applying MCD to each layer in $\mathbb{L}$ according to Eqs. (3) and (4).
7:         $g_{s}^{t}\leftarrow\nabla_{x_{\mathrm{adv}}^{t-1}}\mathcal{L}\left(f_{\mathcal{M}}(\mathcal{T}(x_{\mathrm{adv}}^{t-1}))\right)$
8:     end for
9:     $g^{t}\leftarrow\frac{1}{S}\sum g_{s}^{t}$
10:     Update $x_{\mathrm{adv}}^{t}$ with gradient $g^{t}$ using iterative methods (like MI-FGSM [7]).
11:     Project ${x_{\mathrm{adv}}^{t}}$ into the $\epsilon$-ball of $x$.
12:end for
13:return $x_{\mathrm{adv}}^{T}$
Unlike standard Dropout or DropConnect in training neural networks, MCD does not need additional re-scaling. The random masks $\mathcal{M}_{w}$ and $\mathcal{M}_{b}$ in Eq. 3 are re-generated for each inference, producing diverse variants of the surrogate model with different parameters if we conduct multiple inferences at an iteration. In addition, MCD  can be naturally integrated with existing input transformation and gradient stabilization methods for crafting adversarial examples, as shown in Lines 5 and 8 in Algorithm 1.
4.3 Discussion
In this section, we compare MCD with related self-ensemble methods in more details.
While MCD is proposed from the perspective of mitigating perturbation co-adaptation, it turns out to be a self-ensemble method that constructs multiple new models at each iteration. Existing self-ensemble method [18] targets object detection task, which is different from ours. SE-ViT [28] is specifically designed for vision Transformer surrogate models; as shown in Sec. 5, its ASR is lower than MCD even when using ViT as surrogate model. More closely related are Ghost Network [22] and MUP [45], but are much outperformed by MCD (c.f. Sec. B.2 and Sec. 5.2).
We now analyze the effectiveness of MCD from model ensemble perspective. It has been hypothesized that an adversarial image that remains adversarial for multiple models is more likely to transfer to other models [26]. MCD generates independent random masks for each selected layer and also at each optimization iteration. In this sense, it brings in more randomness and further diversification than Ghost Network and MUP. Specifically, Ghost Network perturbs only skip connections for residual networks (like ResNet-50), while MUP masks unimportant parameters according to a predefined metric at each iteration. This observation is also in accordance with [18, 3, 25], which show that increasing the number of surrogate models generally enhances the transfer attack performance. On the other hand, from the perspective of ensemble techniques in machine learning, each variant model should also be informative about or useful to the targeted task of image classification [18, 3]. As empirically shown in Sec. 5.5, MCD achieves a good tradeoff in terms of model diversity and utility, thereby enhancing the overall attack performance.
In addition, MCD is simple to implement and is robust across a range of parameter choices. In particular, our implementation of MCD is to pick the same DropConnect probability, i.e., $p_{w}=p_{e}=p$, for selected layers. MCD can be applied to all layers in the model with little decrease of ASR, compared to our best practice that only selects the linear and normalization layers in this paper. See the experiments in Sec. 5 for more details.
5 Experiments
This section conducts extensive experiments to show the effectiveness of the proposed method. In particular, we consider both CNN- and Transformer-based classifiers.
5.1 Experimental Settings
Dataset
We utilize the ImageNet-compatible dataset,https://github.com/cleverhans-lab/cleverhans/tree/master/cleverhans_v3.1.0/examples/nips17_adversarial_competition/dataset which served as the official dataset for the NIPS 2017 Attack Challenge. This dataset contains both ground-truth and targeted labels, making it well-suited for conducting experiments on targeted attacks.
General Setting
We adopt the $\ell_{\infty}$-norm as the constraint on perturbation, with budget $\epsilon=16/255$. The learning rate is chosen as $\alpha=2/255$. We use the Logit loss [47] as our objective function in Eq. 1. By default, we set the maximum number of optimization iterations to $1,000$ and the batch size to $32$ for all baseline methods, to ensure sufficient optimization. Furthermore, we notice that different attack methods may take different amounts of computations per iteration. We hence fix the same number of inferences for each optimization iteration to maintain fair comparisons across all attack methods.
Source : Incv3
Target model
Attack RN18 RN50 VGG16 Incv3 EFB0 DN121 MBv2 IRv2 Incv4 Xcep Avg.
DI 2.2 3.9 3.4 99.1 3.6 5.0 1.2 7.7 8.9 7.0 14.2
RDI 5.8 5.5 3.9 99.0 8.0 8.5 3.8 18.6 18.8 11.1 18.3
SI 6.7 6.7 4.3 98.8 9.7 9.7 4.4 23.3 22.1 13.6 19.9
MUP 13.9 13.6 9.6 98.4 17.7 22.4 8.1 42.2 42.2 26.4 29.5
BSR 15.8 13.9 11.9 98.7 20.5 24.3 9.6 45.7 45.5 30.3 31.6
Admix 18.5 16.7 13.8 98.1 23.8 27.5 15.9 46.3 47.1 38.9 34.7
ODI 14.4 22.3 22.0 99.4 26.0 39.5 13.9 51.8 60.7 44.7 39.5
CFM 37.4 37.9 27.3 97.9 46.1 53.0 27.8 76.9 76.1 68.2 54.9
MCD 51.9 52.5 42.1 97.5 60.6 68.6 48.7 87.1 88.1 84.1 68.1
Source : ViT
Target model
Attack RN18 RN50 VGG16 Incv3 EFB0 DN121 MBv2 IRv2 Incv4 Xcep Avg.
DI 0.5 1.0 0.8 2.1 2.1 1.1 1.3 1.8 1.6 1.4 1.4
RDI 1.7 2.4 1.3 4.1 6.3 4.1 2.7 5.8 4.9 4.6 3.8
SI 2.9 3.9 1.1 8.0 9.2 5.9 2.7 7.9 6.2 6.8 5.5
BSR 5.0 8.2 3.9 11.2 15.0 12.8 5.3 15.5 13.4 11.2 10.2
SE-ViT 8.9 12.0 6.9 21 25.3 20.5 8.3 23.8 22.5 20.3 17.0
Admix 16.4 20.4 13.2 31.8 38.3 31.5 17.8 35.3 31.9 29.8 26.6
ODI 12.4 20.0 10.6 28.3 28.9 30.9 10.4 35.5 34.4 27.9 23.9
CFM 26.1 33.4 18.0 45.2 56.8 47.3 23.2 54.5 49.9 46.2 40.1
MCD 36.2 43.9 28.2 58.0 69.0 56.7 35.8 63.2 62.9 57.8 51.2
Table 1: ASRs (%) against ten target models on the ImageNet-Compatible dataset. All the attack methods are combined with MI-TI. The best results are shown in bold and the second best results are underlined.
Surrogate and Target Models
Our experiments choose various models commonly used in the literature [2]. These include 1) CNN-based models: VGG-16 [32], ResNet-18 (RN18) [15], ResNet-50 (RN50) [15], DenseNet-121 (DN121) [17], Xception (Xcep) [4], MobileNet-v2 (MBv2) [31], EfficientNet-B0 (EFB0) [37], Inception ResNetv2 (IRv2) [36], Inception-v3 (Incv3) [35], and Inception-v4 (Incv4) [36]; and 2) Transformer-based models: ViT [9], LeViT [12], ConViT [10], Twins [5], and Pooling-based Vision Transformer (PiT) [16]. All the models are pre-trained on the ImageNet dataset [6] . Additionally, we include CLIP [29], trained on 400 million text-image pairs, to evaluate the transferability across different modalities. When using Transformer-based models, the input image is resized to 224$\times$224 to meet the model input requirement.
Baseline Attack Methods
We compare MCD with various existing transfer-based methods, including DI [44], RDI [48], SI [24], Admix [42], BSR [40], ODI [1], and CFM [2]. We also include two existing self-ensemble methods, namely, MUP (whose implementation only handles CNN layers) [45] and SE-ViT (which is specifically designed for vision Transformers) [28]. These methods are primarily used in combination with TI-FGSM [8] and MI-FGSM [7] during the optimization process. It is worth noting that some of these methods, namely, SI, BSR, Admix, CFM, MUP and SE-ViT, are implemented together with RDI, which have been reported to obtain higher transfer ASRs [2]. As previously mentioned, the attack methods are configured with an identical number of inferences per iteration, denoted as $S$. Specifically, we pick $S$ scaled copies in SI and in the inner loop of Admix, and $S$ transformed images for BSR . For other baselines, we perform $S$ forward-backward passes and use the average gradient on $x_{\mathrm{adv}}$ to update the adversarial example in each iteration. In the main experiment, we will set $S=5$; comparison of other choices is studied in Sec. 5.3.
MCD Setting
While the DropConnect probabilities can be chosen differently for the weight and bias parameters in a linear layer (or the transformation parameters in the normalization layer) and also across different layers, we simply pick the same probability for all selected parameters, that is, $p_{w}=p_{b}=p$, which greatly simplifies the implementation. Through our ablation study, we find that MCD performs well across a range of probabilities. For our experiments, we will select the following DropConnect probabilities: $0.05$ for ResNet-50, $0.02$ for Inception-v3, $0.04$ for DenseNet-121, $0.01$ for Vision Transformer, and $0.03$ for CLIP. By default, MCD applies DropConnect to all linear and normalization layers in the surrogate model.
Source: RN50 Source: DN121
Attack ViT LeViT ConViT Twins PiT CLIP Avg. ViT LeViT ConViT Twins PiT CLIP Avg.
DI 0.4 6.7 0.6 3.8 1.8 0.5 2.3 0.3 4.0 1.2 2.0 2.1 0.3 1.7
RDI 2.8 24.0 4.4 12.6 10.1 1.2 9.2 1.0 12.0 2.1 6.9 8.2 1.4 5.3
SI 8.0 42.9 7.7 25.1 23.3 3.7 18.4 3.9 21.5 4.4 10.2 14.3 2.1 9.4
BSR 6.8 42.4 7.7 25.3 21.9 2.3 17.7 2.7 21.6 2.7 10.6 11.9 1.2 8.5
Admix 7.2 43.4 5.9 22.7 19.4 4.2 17.1 4.2 33.8 4.0 18.0 19.1 3.9 13.8
ODI 15.5 49.3 11.8 31.6 35.1 5.6 24.8 8.5 36.3 8.4 19.8 26.7 4.9 17.4
MUP 7.0 48.3 8.2 30.7 26.5 3.6 20.7 5.2 38.9 5.0 23.2 20.1 4.1 16.1
CFM 17.3 65.8 14.6 47.5 39.9 7.9 32.2 10.4 49.0 8.3 30.1 30.4 5.5 22.3
MCD 34.5 76.7 27.1 60.1 57.1 15.7 45.2 28.0 68.5 23.2 53.8 53.6 14.1 40.2
Table 2: ASRs (%) against five transformer-based target models on the ImageNet-Compatible dataset. We also include the result on CLIP. All the attack methods are combined with MI-TI. The best results are shown in bold and the second best results are underlined.
(a) $S=1$
(b) $S=5$
(c) $S=10$
Figure 3: Average ASRs along optimization iterations. Here $S$ denotes the number of inferences per iteration.
5.2 Main Result
We first study the performance of MCD on the ImageNet-Compatible dataset. We employ ResNet-50, Inception-v3, DenseNet-121, and ViT as surrogate models, and evaluate the obtained adversarial examples on 16 target models.
Tab. 1 reports the ASRs of various attack methods on ten CNN-based target models. MCD achieves the best average ASR. Particularly, with Inception-v3 as surrogate model, the ASRs are increased by 14.8% and 20.9% for the challenging target models VGG16 and MBv2, respectively. When transferring from Transformer-based model ViT to CNN-based models, MCD again attains the best average ASR of 51.2%. We also report the results of self-ensemble methods MUP [45] and SE-ViT [28], and MCD clearly outperforms these two methods by a large margin.
In Tab. 2, we present experimental results of targeted attacks, where adversarial examples are generated using CNN-based models and are transferred to Transformer-based neural networks. This task is considered more challenging in the context of transfer-based attacks [27], as the ASRs in this case are still relatively low. Again, our method significantly improves the attack performance over existing methods: it increases the average ASR from 32.2% to 45.2% with ResNet-50 as surrogate model, and from 22.3% to 40.2% when the surrogate model is DenseNet-121.
Due to space limit, additional experimental results with different surrogate models can be found in Sec. B.4. We also visualize the heatmaps of some generated adversarial examples in Appendix A for qualitative comparison.
5.3 More Iterations and Inferences
We now study the performance given different total iterations and different numbers of inferences per iteration, which are denoted as $T$ and $S$, respectively. We use ResNet-50 as the surrogate model and evaluate how well the adversarial examples are transferred to the 16 target models.
Figure 4: Average ASRs with different iterations (denoted by $T$) and different numbers of inferences per iteration (denoted by $S$).
The results are reported in Fig. 3. Although existing methods may also benefit from additional optimization iterations, MCD and Admix have the best gains when $T$ increases, while MCD achieves a much higher ASR than Admix. With $S$ increasing, MCD can outperform CFM even at an early stage of the optimization process. We also depict Fig. 4 to ease the comparison of the gains of different methods when both $T$ and $S$ increase. As we observe, MCD benefits the most from an additional compute budget.
5.4 Ablation Study
In this section, we conduct an ablation study to investigate the impacts of 1) different types of layers where DropConnect is applied, and 2) different DropConnect probabilities.
Source Model BN FC Conv ASRs (%)
RN50 72.1
41.1
65.1
72.4
69.1
64.7
69.2
Source Model LN FC - ASRs (%)
ViT 58.6
63.9
65.2
Table 3: Applying DropConnect to different types of layers. The reported ASRs(%) are averages over 16 different target models.
Different Types of Layers
We use ResNet-50 and ViT as surrogate models to analyze the impacts of layer types. For ResNet-50, we apply MCD to different combinations of Batch Normalization (BN) layer, Fully Connected (FC) layer, and Convolutional (Conv) layer. Similarly, we consider Layer Normalization (LN) and FC layers (including linear transformation layer in the attention layer) for ViT.
Tab. 3 presents the experimental results. Notably, simply applying MCD to all layers achieves equal or higher ASRs, compared with other baselines. The combination of BN (or LN) and FC layers performs the best, which validates our implementation in Sec. 4. For ResNet-50, applying DropConnect to Conv layers performs worse than BN layers. We conjecture that Conv layers generally have sparser weights and may be less affected by perturbation co-adaptation.
DropConnect Probability
We further investigate the impact of varying DropConnect probabilities, denoted as $p$, using ResNet-50 as a surrogate model. We run MCD with different probabilities ranging from $0.01$ to $0.09$ and report the average ASRs over 16 different models. Along different $p$, the mean of average ASRs is 66.3% with a standard deviation of 5.9%. The highest average ASR is 72.4%, achieved at $p=0.05$. Notably, with $p\in[0.03,0.07]$, MCD consistently outperforms other baselines by more than 2%, underscoring the stability of the proposed method across different choices. Due to space limit, the detailed results are given in Fig. 9 in the supplementary material.
5.5 Model Diversity and Utility
In this section, we show the diversity and utility of constructed models. Here diversity is measured by first computing the mean KL-divergence between the output distributions of each pair of model variants for an input and then averaging these values over the ImageNet-compatible dataset. Similarly, the utility is quantified by the mean top-1 accuracy over the model variants and the dataset. We use the same surrogate model ResNet-50 as in the ablation study.
The results are presented in Fig. 5. As DropConnect probability $p$ increases, the diversity among different variants improves while the utility decreases. As shown in Fig. 9 in the supplementary material, the average ASR using the same surrogate model increases with $p\in[0.01,0.05]$ and then decreases with $p\in[0.05,0.09]$, while the highest ASR is obtained at $p=0.05$. This verifies our discussion in Sec. 4.3 that the attack performance is affected by both diversity and utility. We conclude that the better performance of MCD is due to the improved diversification across the variants while keeping each variant sufficiently useful.
5.6 Attack Performance Against Defense Methods
Finally, we evaluate MCD against several defense mechanisms, including adversarial training methods adversarially trained ResNet-50 (advRN) [30] and Ensemble-Adversarial-Inception-ResNet-v2 (ensIR) [19], and also High-level representation Guided Denoiser (HGD) [23], Bit Depth Reduction (Bit) [14], JPEG compression [14], and R&P [43]. We utilize ResNet-50 as the surrogate model. For Bit, JPEG, and R&P, the target model is ResNet-18. As shown in Tab. 4, MCD  outperforms all other baselines. Notably, in the challenging scenario of ensIR and HGD defenses, our method surpasses the second best ASRs by 29.4% and 10.5%, respectively.
(a) Diversity
(b) Utility
Figure 5: ASRs(%) with increasing mask ratio.
RN50 Defense methods
Method advRN ensIR JPEG Bit R&P HGD
DI 10.6 0.0 26.7 50.5 41.1 0.1
RDI 39.6 0.8 58.8 75.8 71.3 0.8
SI 61.8 9.4 75.1 80.7 78.8 2.1
Admix 68.9 5.7 76.9 81.4 78.6 2.4
BSR 60.4 3.0 76.8 85.2 83.4 2.6
ODI 58.6 5.1 71.7 73.8 76.1 0.3
CFM 84.1 13.8 88.2 91.5 89.3 15.2
MCD 88.2 43.2 91.2 92.2 92.7 25.7
Table 4: ASRs (%) against six defense methods.
6 Concluding Remarks
In this paper, we first show the perturbation co-adaptation issue in existing transfer-based attack methods, and then propose MCD that creates diverse variants of surrogate model at each iteration to improve transferability. MCD brings in more diversification across the variants while maintaining the utility of each variant. Extensive experiments demonstrate the superior performance of MCD.
Supplementary Material
Appendix A Visualization of Adversarial Examples
Figs. 6 and 7 visualize the attention heatmap of the adversarial examples, with surrogate model ResNet-50 and target model ResNet-18. We observe that the attention heatmaps w.r.t. MCD on the surrogate model are more similar to the heatmaps on target model, compared with other attack methods. This in part explains that the generated adversarial example of MCD is more transferable to the target model.
Figure 6: Attention heatmap of adversarial example. The true label is ‘black swan’ and the target label is ‘weasel’. The intensity of red indicates the level of importance assigned to each area, influencing the classifier prediction towards the target label.
Figure 7: Attention heatmap of adversarial example. The true label is ‘cinema’ and the target label is ‘croquet ball’. The intensity of red indicates the level of importance assigned to each area, influencing the classifier prediction towards the target label.
Appendix B Additional Experimental Results
B.1 Additional Results in Pilot Study
We present more pilot study results discussed in Sec. 4.1.
(a) DenseNet-121 $\to$ ResNet-18
(b) DenseNet-121 $\to$ MobileNet-v2
Figure 8: ASRs(%) with increasing mask ratio.
B.2 Initial Results of Applying Dropout to Perturbation and Intermediate Features
Tab. 5 reports the results of applying random masking to perturbation as well as to intermediate features. We see that the Dropout idea generally helps with increased ASR, but does not achieve the best performance. Notice that CFM is also implemented in combination with RDI, which we write it explicitly as RDI+CFM in the table.
We also include the empirical result of Ghost Networks (GN) [22]. It perturbs only skip connections for residual networks and also uses one new model at each iteration. For a fair comparison, we combine it with RDI and pick a larger iteration number for optimization. However, its performance is still much lower. Compared with RDI+DropF, we can see that the additional randomness of applying Dropout to each layer is beneficial to transfer-based targeted attack.
Method DI GN DI+DropP DI+DropF RDI RDI+DropP RDI+DropF RDI+CFM
ASR(%) 26.8 29.7 38.7 60.0 42.0 48.7 63.0 66.1
Table 5: Average ASRs of applying Dropout onto perturbation and intermediate features, denoted as +DropP and +DropF, respectively. Results are the average over 16 target models on the ImageNet-compatible dataset, with ResNet-50 as surrogate model. Detailed experimental setting can be found in Section 5.1.
B.3 Ablation Study on DropConnect Probability
Experimental results of varying DropConnect probabilities are shown in Fig. 9, indicating a stable performance of MCD across a range of different parameter values.
Figure 9: ASRs(%) averaged over 16 different models with varying DropConnect probabilities, using ResNet-50 as the surrogate model.
B.4 Additional Attack Results
Tab. 6 and Tab. 7 present additional attack results. The setting is the same as described in Sec. 5.1. Our method MCD outperforms other baselines by a large margin. Furthermore, in the setting where CLIP is used as a surrogate model, MCD again results in a notable improvement. Yet, the ASRs are still relatively low, indicating that there is much room for further enhancement, particularly in the case of transferring from CLIP to CNN-based models.
Source : RN50
Target model
Attack RN18 RN50 VGG16 Incv3 EFB0 DN121 MBv2 IRv2 Incv4 Xcep Avg.
DI 60.1 98.7 64.6 8.1 28.6 77.3 28.4 13.5 21.1 15.1 41.5
RDI 80.6 98.7 75.4 32.8 52.6 87.4 49.6 47.8 51.1 40.7 61.7
SI 82.9 98.8 73.7 60.8 66.0 90.0 59.4 66.9 68.7 60.6 72.8
BSR 87.9 98.5 84.1 54.9 70.0 91.6 65.3 66.5 68.9 63.3 75.1
Admix 83.2 98.1 78.1 57.1 67.8 86.8 70.3 61.9 64.4 64.5 73.2
ODI 77.5 98.9 83.0 63.9 69.9 89.8 61.2 71.0 71.1 69.2 75.5
MUP 90.3 97.7 85.6 61.8 75.0 91.6 79.3 70.9 71.8 68.7 79.3
CFM 92.0 98.0 89.5 74.8 85.6 93.4 85.2 82.6 82.5 81.7 86.5
MCD 93.0 98.2 91.0 82.6 88.5 93.6 90.9 82.1 85.5 85.0 89.0
Source : DN121
Target model
Attack RN18 RN50 VGG16 Incv3 EFB0 DN121 MBv2 IRv2 Incv4 Xcep Avg.
DI 28.1 36.4 35.8 6.5 14.5 99.0 10.1 9.8 13.9 10.1 26.4
RDI 50.9 54.4 46.7 19.3 26.9 98.6 20.6 29.8 30.5 24.7 40.2
SI 54.8 60.9 43.0 35.6 37.9 98.6 26.7 43.6 43.5 34.6 47.9
BSR 61.6 70.3 57.3 34.1 41.7 98.3 30.2 42.7 44.5 37.5 51.8
Admix 72.4 74.3 65.8 46.9 55.1 97.9 51.2 55.5 56.6 53.9 63.0
ODI 59.6 71.5 69.6 48.1 52.2 98.9 35.5 58.1 60.0 52.7 60.6
MUP 79.9 86.9 75.2 50.3 60.5 97.7 51.4 65.7 66.5 61.0 69.5
CFM 82.4 88.9 78.6 60.9 69.9 98.5 60.5 73.0 72.7 70.3 75.6
MCD 89.6 90.2 85.1 77.8 84.4 97.4 84.4 83.2 84.1 83.7 86.0
Source : CLIP
Target model
Attack RN18 RN50 VGG16 Incv3 EFB0 DN121 MBv2 IRv2 Incv4 Xcep Avg.
DI 0.1 0.1 0.1 0.0 0.0 0.1 0.2 0.0 0.0 0.1 0.1
RDI 0.5 0.4 0.2 0.4 0.9 0.4 0.3 0.4 0.2 0.7 0.4
SI 0.6 0.5 0.6 1.4 1.6 1.4 1.0 1.0 0.8 1.2 1.0
BSR 1.7 1.6 0.9 1.9 3.5 2.4 1.5 1.9 2.3 2.2 2.0
Admix 2.7 2.2 1.6 4.1 6.0 3.5 2.8 2.9 3.8 3.6 3.3
ODI 7.7 7.0 5.1 11.0 13.7 13.0 6.5 11.3 11.8 11.5 9.9
CFM 3.1 3.3 1.4 4.9 7.8 4.3 2.5 5.4 4.1 4.8 4.2
MCD 8.8 9.5 4.9 14.6 17.5 14.9 8.5 12.5 12.8 12.4 11.6
Table 6: Additional experimental results against ten target models on the ImageNet-Compatible dataset, using RN50, DN121 and CLIP as surrogate model, respectively. All the attack methods are combined with MI-TI. The best results are shown in bold and the second best results are underlined.
Source : Incv3
Target model
Attack ViT LeViT ConViT Twins PiT CLIP Avg.
DI 0.1 0.8 0.1 0.3 0.5 0.2 0.3
RDI 0.1 3.1 0.6 2.0 1.9 0.0 1.3
SI 1.0 5.0 1.0 2.0 3.8 0.4 2.2
BSR 1.3 15.3 1.9 5.6 7.3 0.5 5.3
Admix 2.0 17.4 1.9 6.7 8.5 1.4 6.3
ODI 3.6 21.6 4.4 8.9 15.1 2.0 9.3
MUP 1.4 11.9 2.3 4.4 5.4 0.4 4.3
CFM 9.7 43.4 8.8 23.2 26.4 3.9 19.2
MCD 14.7 57.6 16.7 34.5 38.0 5.3 27.8
Source : ViT
Target model
Attack ViT LeViT ConViT Twins PiT CLIP Avg.
DI 100.0 11.3 15.6 7.4 14.8 3.2 25.4
RDI 100.0 31.5 38.9 23.6 36.2 7.7 39.6
SI 100.0 48.2 59.2 28.3 51.4 14.8 50.3
BSR 100.0 57.8 63.0 45.7 67.2 18.5 58.7
Admix 99.8 78.2 81.5 66.4 79.2 39.6 74.1
SE 100 65.6 75.7 56.4 72 20.4 65.0
ODI 100.0 74.8 68.3 63.1 81.7 31.6 69.9
CFM 99.9 90.9 94.2 82.5 91.6 55.3 85.7
MCD 99.6 92.4 95.7 87.4 94.4 63.5 88.8
Source : CLIP
Target model
Attack ViT LeViT ConViT Twins PiT CLIP Avg.
DI 0.0 0.0 0.1 0.0 0.0 99.8 16.6
RDI 1.3 1.5 0.6 0.4 1.0 99.8 17.4
SI 3.1 4.5 2.5 1.7 6.2 99.6 19.6
BSR 7.0 9.2 4.9 2.9 8.5 99.5 22.0
Admix 7.3 11.9 5.5 4.2 11.1 99.0 23.2
ODI 19.1 24.3 13.1 11.7 22.7 99.8 31.8
CFM 13.3 15.4 8.6 4.7 13.5 99.8 25.9
MCD 32.5 35.5 25.1 16.7 33.0 98.4 40.2
Table 7: Additional experimental results against five transformer-based target models on the ImageNet-Compatible dataset, using RN50, DN121 and CLIP as surrogate model, respectively. All methods are combined with MI-TI. The best results are shown in bold, and the second best results are underlined.

### v2

@@
-A Simple DropConnect Approach to Transfer-based Targeted Attack
+RaPA: Enhancing Transferable Targeted Attacks via Random Parameter Pruning
 Abstract
-We study the problem of transfer-based black-box attack, where adversarial samples generated using a single surrogate model are directly applied to target models.
-Compared with untargeted attacks, existing methods still have lower Attack Success Rates (ASRs) in the targeted setting, i.e., the obtained adversarial examples often overfit the surrogate model but fail to mislead other models.
-In this paper, we hypothesize that the pixels or features in these adversarial examples collaborate in a highly dependent manner to maximize the success of an adversarial attack on the surrogate model, which we refer to as perturbation co-adaptation.
-Then, we propose to Mitigate perturbation Co-adaptation by DropConnect (MCD) to enhance transferability, by creating diverse variants of surrogate model at each optimization iteration.
-We conduct extensive experiments across various CNN- and Transformer-based models to demonstrate the effectiveness of MCD.
-In the challenging scenario of transferring from a CNN-based model to Transformer-based models, MCD achieves 13% higher average ASRs compared with state-of-the-art baselines.
-MCD  boosts the performance of self-ensemble methods by bringing in more diversification across the variants while reserving sufficient semantic information for each variant.
-In addition, MCD attains the highest performance gain when scaling the compute of crafting adversarial examples.
-footnotetext: Preprint. Work in progress.
+Compared to untargeted attacks, targeted transfer-based attack is still suffering from much lower Attack Success Rates (ASRs), although significant improvements have been achieved by kinds of methods, such as diversifying input, stabilizing the gradient, and re-training surrogate models.
+In this paper, we find that adversarial examples generated by existing methods rely heavily on a small subset of surrogate model parameters, which in turn limits their transferability to unseen target models.
+Inspired by this, we propose the Random Parameter Pruning Attack (RaPA), which introduces parameter-level randomization during the attack process.
+At each optimization step, RaPA randomly prunes model parameters to generate diverse yet semantically consistent surrogate variants.We show this parameter-level randomization is equivalent to adding an importance-equalization regularizer, thereby alleviating the over-reliance issue.
+Extensive experiments across both CNN and Transformer architectures demonstrate that RaPA substantially enhances transferability.
+In the challenging case of transferring from CNN-based to Transformer-based models, RaPA achieves up to 11.7% higher average ASRs than state-of-the-art baselines(with 33.3% ASRs), while being training-free, cross-architecture efficient, and easily integrated into existing attack frameworks.
 1 Introduction
-In recent years, deep neural networks have become prevalent in computer vision applications [15, 17, 9], but are highly vulnerable to maliciously crafted inputs, which are called adversarial examples [34, 11].
-A major concern is their transferability, that is, adversarial examples generated using a white-box model can be directly used to fool other black-box models, without any access to their architectures, parameters or gradients [13].
+Deep neural networks have become prevalent in computer vision applications [16, 18, 10], but are highly vulnerable to maliciously crafted inputs, called adversarial examples [41, 12].
+A major concern is their transferability, that is, adversarial examples generated using a white-box model can directly fool other black-box models, without any access to their architectures, parameters or gradients [14].
 Since this type of attacks, usually referred to as transfer-based attacks, do not require any interaction with the target model, they pose severe security risks to real-world machine learning systems.
-Therefore, studying effective transfer-based attack methods is crucial to understanding the vulnerabilities, and in turn enhancing the robustness of machine learning models.
+Therefore, studying effective transfer-based attack methods is crucial to understand the vulnerabilities and further enhance model robustness.
 This paper focuses on targeted transfer-based attacks with a single surrogate model, where the goal is to deceive black-box models to classify input images into a specific incorrect category.
@@
 A key observation is that the generated adversarial examples tend to overfit the surrogate model but fail to generalize to other models.
-To improve transferability, various strategies have since been proposed.
-Noticing that multiple surrogate models can help enhance transferability but in practice finding proper models for the same task is not easy [7, 26, 25, 46], model self-ensemble [18, 45, 22, 28] tries to create multiple models from only one accessible model, which is in accordance with our setting.
-Input transformation [44, 48, 1, 42, 40, 2] applies different transformations to inputs and diversify input patterns in order to reduce overfitting.
-A notable method is Clean Feature Mixup (CFM) [2], which randomly mixes high-level features with shuffled clean features and has achieved state-of-the-art performance.
+To improve transferability, various strategies have been proposed.
+Observing that multiple surrogate models can help enhance transferability but in practice finding proper models for the same task is not easy [8, 30, 29, 55], model self-ensemble [19, 54, 25, 33] tries to create multiple models from an accessible model.
+Input transformation [52, 57, 1, 48, 46, 49, 2, 26] applies different transformations to inputs and diversify input patterns to reduce overfitting.
+A notable method is Clean Feature Mixup (CFM) [2], which randomly mixes high-level features with shuffled clean features.
+Building upon it, Feature Tuning Mixup (FTM) [26] introduces learnable and attack-specific feature perturbations, achieving new state-of-the-art performance in transferability.
 Despite these progresses, there is still much room for further improvement.
-In this work, we hypothesize that there is a key aspect underlying the present overfitting issue: the pixels in many adversarial examples, obtained from existing transferred-based methods, collaborate in a highly dependent manner to maximize the success of adversarial attacks on the surrogate model, which we refer to as perturbation co-adaptation.
-Similar to the co-adaption issue of units in training neural networks [33], we believe that perturbation co-adaptation would lead to overly complex patterns that can fool the surrogate model but fail on other target models.
-How to further alleviate this co-adaptation issue would be a key to improving the transferability of adversarial examples over existing transfer-based attack methods.
-To this end, we propose to Mitigate perturbation Co-adaptation by DropConnect (MCD), motivated by the classic Dropout technique that can effectively prevent units from co-adaptation during training neural networks [39].
-Rather than directly applying Dropout to the input or the perturbation, our method extends random masking to 1) high-level features that contain richer semantic information, and 2) model parameter space for more flexible and diverse feature disturbances for perturbation.
-Fig.
-1 illustrates the proposed method.
-At each optimization iteration, MCD  creates diverse variants of the surrogate model with masked parameters, which are then used to update the adversarial examples.
-In this sense, the final algorithm can be regarded as a self-ensemble method.
-Compared with other related self-ensemble methods [45, 22], MCD brings in more diversification across the variants while reserving sufficient semantic information for each variant, hence improves the transfer attack performance.
-In addition, MCD can be easily implemented into all layers in the surrogate model, with little comprise in terms of attack performance.
-A more concrete comparison is left to Sec.
-4.3.
-Figure 1: An illustration of the proposed method MCD. We apply MCD to selected layers in the surrogate model to create multiple and diverse variants at each optimization iteration.
+In this work, we take a different perspective and identify a previously overlooked cause of poor transferability: the generated adversarial perturbations rely excessively on a small subset of parameters in the surrogate model, which limits their generalization to other models that have different parameter configurations or training dynamics.
+In other words, adversarial perturbation in existing methods tends to exploit a few “shortcut” parameters, leading to strong white-box performance but poor black-box transferability.
+To mitigate this issue, we propose Random Parameter Pruning Attack (RaPA) that introduces parameter-level randomization into the attack process.
+At each optimization step, RaPA randomly prunes a subset of parameters in the surrogate model and uses multiple masked variants to update the adversarial example.
+We show that taking the expectation over such random masks is equivalent to adding an importance regularization term that aims to equalize parameter contributions, thus preventing over-reliance on a few dominant parameters.
+Conceptually, RaPA can be viewed as a self-ensemble method: each randomly pruned model represents a diverse yet semantically consistent variant of the surrogate.
+Previous self-ensemble approaches SASD-WS [50], MUP [54], and Ghost Network [25] rely on training-based model enhancement, deterministic pruning metrics, and structural perturbations, respectively.
+In contrast, RaPA is training-free, cross-architecture efficient, and straightforward to implement.
+Figure 1: An illustration of the proposed method RaPA. We apply RaPA to selected layers in the surrogate model to create multiple and diverse variants at each iteration.
 We evaluate the proposed method across various CNN- and Transformer-based target models, as well as against several defense methods.
-The experimental results show that MCD outperforms other state-of-the-art methods.
-In particular, in the challenging scenario of transferring from CNN-based model to Transformer-based models, MCD achieves 13.0% and 17.9% higher average ASRs with ResNet-50 [15] and DenseNet-121 [17] as surrogate models, respectively.
-Moreover, MCD achieves the highest performance gain when scaling the compute for crafting adversarial examples.
+The experimental results show that RaPA outperforms other state-of-the-art methods.
+In particular, in the challenging scenario of transferring from CNN-based model to Transformer-based models, RaPA achieves 11.7% and 17.5% higher average ASRs with ResNet-50 [16] and DenseNet-121 [18] as surrogate models, respectively.
+Moreover, RaPA achieves the highest performance gain when scaling the compute for crafting adversarial examples.
 Specifically, with ResNet-50 as surrogate model, increasing the optimization iterations from 300 to 500 and number of forward-backward passes per iteration from 1 to 5 boosts the average ASR by 15.9%.
 To summarize, our main contributions are as follows:
-  • We show the perturbation co-adaptation issue in existing transfer-based attack methods in the targeted setting.
-Reducing this co-adaptation issue would in turn enhance transferability of adversarial examples.
-  • We propose MCD to improve the transferability, by creating diverse variants of surrogate model during optimization.
-MCD brings in more diversification across the variants while reserving sufficient semantic information for each variant, and is also simple to implement.
-  • Extensive experiments across various surrogate and target models validate the superior performance of MCD over other methods.
-Moreover, MCD benefits the most when we scale the compute of crafting adversarial examples.
-2 Related Works
-In this section, we provide a brief review of related works focusing on transfer-based attack methods.
+  • We show that adversarial examples from existing transfer-based attacks rely heavily on a tiny subset of parameters in the surrogate model.
+Alleviating this over-reliance can in turn enhance the transferability of attack.
+  • We propose the RaPA, which introduces parameter-level randomization during attack optimization.
+We show, both intuitively and empirically, that random pruning implicitly equalizes parameter importance, acting as a regularizer to mitigate the over-reliance issue.
+  • Experiments across diverse surrogate and target models demonstrate that RaPA consistently outperforms existing methods.
+RaPA further benefits from increased computational budget, achieving larger improvements when scaling optimization iterations or inference steps.
+Method DI RDI SI Admix ODI BSR CFM RaPA
+No pruning 98.2 98.7 98.7 98.1 98.9 98.5 98.0 98.2
+Pruning bottom 0.5% 98.2 98.7 98.7 98.0 98.9 98.5 98.0 98.0
+Pruning top 0.5% 16.0 28.6 31.9 29.9 19.9 36.7 51.3 64.5
+Table 1: ASRs (%) before and after pruning the selected subsets of parameters in the surrogate model on the ImageNet-compatible dataset. Detailed experimental setting can be found in Section 4.1.
+2 Preliminary
+This section introduces the background of adversarial attacks and briefly reviews related works on targeted transfer-based attacks. See Appendix A for more related work.
+2.1 Background
+Consider a classification task where the model is defined as a function $f:\mathbb{R}^{n}\to\mathcal{Y}$ that maps an input $x\in\mathbb{R}^{n}$ to a label in the set $\mathcal{Y}$ consisting of all the labels.
+Given a clean image $x$ with its true label $y\in\mathcal{Y}$, untargeted attacks aim to find an adversarial example $x_{\mathrm{adv}}\in\mathbb{R}^{n}$ that is similar to $x$ but misleads the model to produce an incorrect prediction, i.e.
+$f(x_{\mathrm{adv}})\neq y$.
+Here ‘similarity’ is usually measured by an $\ell_{p}$-norm, e.g., $\|x_{\mathrm{adv}}-x\|_{p}\leq\epsilon$ where $\epsilon>0$ is a predefined perturbation budget.
+For targeted attacks, the goal is to modify the model prediction to a particular target label $y_{\mathrm{tar}}$, that is, $f(x_{\mathrm{adv}})=y_{\mathrm{tar}}\neq y$.
+In this work, we will focus on targeted attacks.
+In the white-box setting where the model is fully accessible, adversarial example can be obtained by the following:
+$\operatorname*{arg\,max}_{x_{\mathrm{adv}}}~\mathcal{L}\left(f(x_{\mathrm{adv}})\right),~~\text{s.t.}~~\|x_{\mathrm{adv}}-x\|_{p}\leq\epsilon,$ (1)
+where $\mathcal{L}(\cdot)$ represents a loss function like cross-entropy loss.
+The Fast Gradient Sign Method (FGSM) [12] uses the gradient direction to solve this problem and craft adversarial examples, while Iterative FGSM (I-FGSM) [22] extends this idea to an iterative scheme.
+In particular, at each iteration $t$, the adversarial example is updated by adding a small perturbation:
+$x_{\mathrm{adv}}^{t}=x_{\mathrm{adv}}^{t-1}+\alpha\cdot\mathrm{sign}\left(\nabla_{x_{\mathrm{adv}}^{t-1}}\mathcal{L}\left(f(x_{\mathrm{adv}}^{t-1})\right)\right),$ (2)
+where $x_{\mathrm{adv}}^{0}=x$ and $\alpha>0$ is a step size.
+To make the generated adversarial example satisfy the perturbation budget constraint, a straightforward way is to project $x_{\mathrm{adv}}^{t}$ into the $\epsilon$-ball of $x$.
+2.2 Related Work
+In black-box settings, the gradient information is not available and we only have limited access to the target model.
+Transfer-based method assumes that the adversarial examples generated on one model to mislead not only that model but also other models [41].
+To improve transferability, many methods have been proposed.
+The first class, input-transformation techniques, applies a transformation $\mathcal{T}$ and uses $\nabla_{x_{\mathrm{adv}}^{t-1}}\mathcal{L}\left(f(\mathcal{T}(x_{\mathrm{adv}}^{t-1}))\right)$ as the gradient.
+Diverse Inputs (DI) [52] and its variant Resized DI (RDI) [57] apply random transformations to increase input variation during optimization.
+Translation-Invariant (TI) [9] averages gradients over translated inputs to reduce location sensitivity.
+Structure Invariant Attack (SIA) [49] and Block Shuffle and Rotate (BSR) [46] both perform block-level local transformations, with SIA applying diverse transforms and BSR focusing shuffle and rotate operations.
+Object-based DI (ODI) [1] generates adversarial examples rendered on 3D objects, while Admix [48] mixes inputs with random samples from other classes.
+CFM [2] extends it to the feature space with competing noises, while FTM [26] further adds learnable, attack-specific perturbations to achieve state-of-the-art transferability.
+The second class focuses on stabilizing gradient updates to improve transferability.
+Momentum Iterative FGSM (MI-FGSM) [8] incorporates a momentum term into I-FGSM to help avoid local optima.
+Scale-Invariant (SI) optimization [28] improves transferability by applying perturbations across multiple scaled copies of the input, leveraging the scale-invariance property of deep models.
+Beyond the above methods, another type of approaches re-train the surrogate model to enhance transferability, e.g., DSM [53] and SASD-WS [50] improve model generalization and transferability through knowledge distillation or sharpness-aware self-distillation.
+Closely related to the present work is self-ensemble [19, 54, 25, 33], which creates multiple models from only one surrogate model.
+The self-ensemble method in [33] specifically considers vision Transformer as surrogate model and is denoted as SE-ViT in this paper.
+Ghost Network [25] perturbs surrogate model to create a set of new models and then samples one model from the set at each iteration.
+Masking Unimportant Parameters (MUP) [54] drops out unimportant parameters according to a predefined Taylor expansion-based metric, while Diversity Weight Purning(DWP) [45] only prunes the parameters with small absolute values.
+However, the inherent differences w.r.t.
+model architecture, parameter setting, and training procedure between the surrogate and target models still limit the effectiveness of transfer-based methods on certain models and datasets.
+In the next section, we show that there is a key aspect that renders the current over-fitting issue of transfer-based methods.
+3 Method
+In this section, we first conduct a pilot study to show a key aspect of the overfitting issue in existing transfer-based methods and then propose a random masking based approach.
+Comparison with related methods is also discussed.
+3.1 Motivation
+We observe that the adversarial perturbations generated by solving Problem (1) tend to rely heavily on a small subset of parameters in the surrogate model.
+These parameters may stem from specific training schemes, datasets, or architectural choices.
+As a result, adversarial examples that strongly depend on these parameters often fail to generalize and mislead other models.
+Even with SoTA transfer-based attack methods, this issue remains a key factor that can lead to the failure of adversarial example transfer.
+To quantify the phenomenon, we conduct a pilot study using the framework of Optimal Brain Damage (OBD) [23, 32], which quantifies the importance of each model parameter from the perspective of sensitivity analysis.
+Specifically, given an adversarial example $x_{\mathrm{adv}}$ and a loss function $\mathcal{L}$.
+Let $\theta$} represent the entire set of model parameters.
+The importance of a parameter $\theta_{i}$ in the surrogate model $f$ is computed as:
+$\mathcal{I}(\theta_{i})=\frac{\partial^{2}\mathcal{L}f(x_{\mathrm{adv}})}{\partial\theta_{i}^{2}}\times\theta_{i}^{2}.$ (3)
+This metric reflects how much the loss would change if a parameter $\theta_{i}$ were removed, and can serve as a proxy for its contribution to the effectiveness of the adversarial example.
+Next, we consider pruning two distinct subsets of the surrogate model’s parameters based on this importance metric: the top 0.5% most important and the bottom 0.5% least important parameters(see Section B.1 for details).
+For each adversarial example, we instantiate the model, prune the selected subset, and evaluate the ASR on the resulting model.
+Table 1 reports the ASRs after pruning the two subsets of model parameters.
+Here we use ResNet50 as the surrogate model and detailed setting can be found in Section 4.1.
+We observe that pruning the most important parameters leads to a drastic drop in ASR—more than 46%, whereas pruning the least important parameters yields negligible impact.
+This observation suggests that adversarial examples generated by existing methods are highly dependent on the most important parameters, validating our observation on the over-reliance issue.
+As such, how to further alleviate this strong dependence on specific parameters would be a key to improving the transferability of adversarial examples over existing transfer-based attack methods.
+3.2 Alleviating Over-reliance via Random Parameter Pruning
+As per the pilot study, a direct approach to improving transferability would be to mask the most important parameters at each optimization step, thereby mitigating the over-dependency on them.
+However, accurately identifying important parameters requires computing second-order derivatives, which is computationally expensive for all parameters.
+Although we can approximate them with first-order terms, masking the most important parameters typically causes the surrogate model’s capacity to degrade rapidly, and the resulting adversarial examples may fail to fool the target model—and even the original surrogate model itself (See Section D.1 for further explanations).
+To address this problem, we propose to apply random parameter pruning to the surrogate model at each optimize step.
+This approach avoids expensive computations while achieving the goal of reducing over-reliance on specific parameters.
+Intuition and Theoretical Explanation
+Our core idea is that randomly pruning parameters at different optimization steps encourages the generated adversarial examples to be less dependent on particular parameter subsets.
+This in turn improves transferability across different target models.
+We define a random binary mask $\mathcal{M}\in\{0,1\}^{|\theta|}$, where each entry is independently sampled from a Bernoulli distribution: $\mathcal{M}_{i}\sim\text{Bernoulli}(1-p)$.
+Here $p\in[0,1]$ is the probability of masking a parameter.
+With a small $p$, we would have $\mathbb{E}[\mathcal{M}_{i}]\approx 1$.
+Then the parameter of the model used in the forward pass becomes $\mathcal{M}\odot\theta$, where $\odot$ denotes element-wise multiplication.
+Under this setup, the expected loss over random masks can be approximated using a second-order Taylor expansion:
+$~\mathbb{E}_{\mathcal{M}}[\mathcal{L}(f(x_{\mathrm{adv}};\mathcal{M}\odot\theta))]$
+$\approx$ $~\mathcal{L}(f(x_{\mathrm{adv}};\theta))+\frac{p(1-p)}{2}\sum_{i}\frac{\partial^{2}\mathcal{L}f(x_{\mathrm{adv}};\theta)}{\partial\theta_{i}^{2}}\theta_{i}^{2},$ (4)
+which is sum of the original loss plus an importance penalty.
+Minimizing this objective while resampling the mask at each step would force the adversarial example to distribute the importance over all parameters, making it more robust to different parameters and thus more transferable.
+Practical Implementation with DropConnect
+The above random parameter pruning method is similar to DropConnect method [44] in training neural networks.
+We notice that DropConnect is mainly effective in terms of linear layers .
+We thus apply DropConnect to the weight and bias parameters of linear layers as well as the transformation parameters of normalization layers.
+Both types of layers are widely used in mainstream architectures including Transformer [43].
+Empirically, our ablation study in Section 4 validates the effectiveness of this choice, compared with convolutional layers.
+We now present our attack method, Random Parameter Pruning Attack (RaPA), as summarized in Algorithm 1.
+Take linear layer for example.
+We perform independent random masking onto the weight and bias (if present) parameters using Bernoulli sampling.
+Specifically, for surrogate model $f$, let $W\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ denote the weight matrix and $b\in\mathbb{R}^{d_{\text{out}}}$ the bias vector associated with a linear layer.
+Here $d_{\text{in}}$ and $d_{\text{out}}$ represent the input and output dimensions, respectively.
+The corresponding masks for the weight matrix and bias are
+$\mathcal{M}_{w}\sim\text{Bernoulli}(1-p_{w}),\mathcal{M}_{b}\sim\text{Bernoulli}(1-p_{b}),$ (5)
+where $\mathcal{M}_{w}\in\{0,1\}^{d_{\text{in}}\times d_{\text{out}}}$ and $\mathcal{M}_{b}\in\{0,1\}^{d_{\text{out}}}$ are the random masks, and $p_{w},p_{b}\in[0,1]$ are DropConnect probabilities.
+Then the masked parameters are computed as
+$W_{\mathcal{M}}=\mathcal{M}_{w}\odot W,~~b_{\mathcal{M}}=\mathcal{M}_{b}\odot b,$ (6)
+where $\odot$ denotes the element-wise multiplication.
+For normalization layer, the same operation is applied similarly to the transformation parameters.
+The random masks are sampled for each selected layer.
+The random masks $\mathcal{M}_{w}$ and $\mathcal{M}_{b}$ in Eq.
+(5) are re-generated for each inference, producing diverse variants of the surrogate model with different parameters if we conduct multiple inferences at an iteration.
+In addition, RaPA can be naturally integrated with existing input transformation and gradient stabilization methods for crafting adversarial examples, as shown in Lines 5 and 8 in Algorithm 1.
+Algorithm 1 Random Parameter Pruning Attack(RaPA)
+1:Classifier $f$; clean image $x$; loss function $\mathcal{L}(\cdot)$; max iterations $T$; $\ell_{p}$ bound $\epsilon$; number of inferences per iteration $S$; DropConnect probabilities $p_{w}$, $p_{b}$; linear and normalization layers $\mathbb{L}$; input transformation $\mathcal{T}$.
+2:Adversarial example $x_{\mathrm{adv}}$
+[…]
+4:for $t=1\to T$ do
+5:  for $s=1\to S$ do
+6:   Obtain modified model $f_{\mathcal{M}}$ by applying RaPA to each layer in $\mathbb{L}$ according to Eqs. (5) and (6).
+7:   $g_{s}^{t}\leftarrow\nabla_{x_{\mathrm{adv}}^{t-1}}\mathcal{L}\left(f_{\mathcal{M}}(\mathcal{T}(x_{\mathrm{adv}}^{t-1}))\right)$
+8:  end for
+9:  $g^{t}\leftarrow\frac{1}{S}\sum g_{s}^{t}$
+10:  Update $x_{\mathrm{adv}}^{t}$ with gradient $g^{t}$ using iterative methods (like MI-FGSM [8]).
+11:  Project $x_{\mathrm{adv}}^{t}$} into the $\epsilon$-ball of $x$.
+12:end for
+13:return $x_{\mathrm{adv}}^{T}$
+Method DI RDI SI Admix ODI BSR CFM RaPA
+All Layer Average 0.32 0.30 0.21 0.12 0.33 0.25 0.19 0.08
+Conv Layer 0.11 0.07 0.03 0.01 0.12 0.05 0.03 0.00
+Norm Layer 0.51 0.52 0.37 0.22 0.53 0.44 0.34 0.15
+Linear Layer 1.00 0.86 0.59 0.18 1.00 0.75 0.55 0.13
+Table 2: Gini coefficients of parameter importance across different layers and methods. Lower values correspond to more uniform parameter importance.
+Analyzing Parameter Importance with Gini Coefficient
+To further verify that our random parameter pruning strategy indeed mitigates the over-reliance on a few dominant parameters, we employ the Gini coefficient to measure the distribution of parameter importance across layers.
+A lower Gini value indicates a more uniform distribution of importance, implying that the adversarial perturbation depends less on specific parameters and generalizes better to unseen models.
+The formal definition and detailed computation process of the Gini coefficient are provided in Sec.
+B.3.
+.
+We compute the Gini coefficients based on the parameter importance values $\mathcal{I}(\theta_{i})$ defined in Eq.
+(3).
+The overall and layer-wise results are summarized in Table 2.
+As shown in Table 2, RaPA achieves the lowest Gini coefficients among all compared methods, suggesting that it effectively flattens the importance distribution and suppresses the over-concentration of sensitivity on a few parameters.
+This balanced importance allocation leads to improved robustness and transferability across different architectures.
+3.3 Discussion
+In this section, we compare RaPA with related self-ensemble methods in more details.
+RaPA was proposed to reduce over-dependence on specific parameters, and it turns out to be a self-ensemble method that constructs multiple new models at each iteration.
+Existing self-ensemble method [19] targets object detection task, which is different from ours.
+SE-ViT [33] is specifically designed for vision Transformer surrogate models; as shown in Section 4, its ASR is lower than RaPA even when using ViT as surrogate model.
+More closely related are Ghost Network [25] and MUP [54], but are much outperformed by RaPA (c.f.
+Section D.1 and Section 4.2).
+We now analyze the effectiveness of RaPA from model ensemble perspective.
+It has been hypothesized that an adversarial image that remains adversarial for multiple models is more likely to transfer to other models [30].
+RaPA generates independent random masks for each selected layer and also at each optimization iteration.
+In this sense, it brings in more randomness and further diversification than Ghost Network, MUP and DWP.
+Specifically, Ghost Network perturbs only skip connections for residual networks (like ResNet-50), while MUP and DWP mask unimportant parameters according to a predefined metricor at each iteration.
+This observation is also in accordance with [19, 3, 29], which show that increasing the number of surrogate models generally enhances the transfer attack performance.
+On the other hand, from the perspective of ensemble techniques in machine learning, each variant model should also be informative about or useful to the targeted task of image classification [19, 3].
+As empirically shown in Section D.4, RaPA achieves a good tradeoff in terms of model diversity and utility, thereby enhancing the attack performance.
+4 Experiments
+Source: RN50 Source: DN121
+Attack ViT LeViT ConViT Twins PiT CLIP Avg. ViT LeViT ConViT Twins PiT CLIP Avg.
+DI 0.4 6.7 0.6 3.8 1.8 0.5 2.3 0.3 4.0 1.2 2.0 2.1 0.3 1.7
+RDI 2.8 24.0 4.4 12.6 10.1 1.2 9.2 1.0 12.0 2.1 6.9 8.2 1.4 5.3
+SI 8.0 42.9 7.7 25.1 23.3 3.7 18.4 3.9 21.5 4.4 10.2 14.3 2.1 9.4
+SIA 3.1 28.4 3.9 16.7 13.5 1.9 11.2 2.4 24.2 2.2 12.9 11.1 1.5 9.0
+BSR 6.8 42.4 7.7 25.3 21.9 2.3 17.7 2.7 21.6 2.7 10.6 11.9 1.2 8.5
+Admix 7.2 43.4 5.9 22.7 19.4 4.2 17.1 4.2 33.8 4.0 18.0 19.1 3.9 13.8
+ODI 15.5 49.3 11.8 31.6 35.1 5.6 24.8 8.5 36.3 8.4 19.8 26.7 4.9 17.4
+MUP 7.0 48.3 8.2 30.7 26.5 3.6 20.7 5.2 38.9 5.0 23.2 20.1 4.1 16.1
+CFM 17.3 65.8 14.6 47.5 39.9 7.9 32.2 10.4 49.0 8.3 30.1 30.4 5.5 22.3
+FTM 18.0 67.6 16.5 47.1 41.5 9.3 33.3 10.7 50.4 9.4 31.1 30.4 4.8 22.8
+RaPA 33.8±1.0 75.4±1.8 27.6±0.2 59.5±0.4 57.3±0.7 15.6±0.1 45.0 27.8±0.1 69.4±0.6 23.5±0.2 53.1±0.6 54.0±0.1 14.1±0.0 40.3
+Table 3: ASRs (%) against five Transformer-based target models on the ImageNet-Compatible dataset.
+All the attack methods are combined with MI-TI.
+The best results are shown in bold and the second best results are underlined.
+This section empirically validates the effectiveness of our method, using both CNN- and Transformer-based models.
+4.1 Experimental Settings
+Dataset
+We utilize the ImageNet-compatible dataset [21], served as the official dataset for the NIPS 2017 Attack Challenge.
+This dataset contains both ground-truth and targeted labels, making it well-suited for targeted-attack.
+General Setting
+We adopt the $\ell_{\infty}$-norm as the constraint on perturbation, with budget $\epsilon=16/255$.
+The learning rate is chosen as $\alpha=2/255$.
+We use the Logit loss [56] as our objective function in Equation 1.
+By default, we set the maximum number of optimization iterations to $1,000$ and the batch size to $32$ for all baseline methods, to ensure sufficient optimization.
+Furthermore, we notice that different attack methods may take different amounts of computations per iteration.
+We hence fix the same number of inferences for each optimization iteration to maintain fair comparisons across all attack methods.
+Surrogate and Target Models
+Our experiments choose various models commonly used in the literature [2].
+These include 1) CNN-based models: VGG-16 [38], ResNet-18 (RN18) [16], ResNet-50 (RN50) [16], DenseNet-121 (DN121) [18], Xception (Xcep) [4], MobileNet-v2 (MBv2) [37], EfficientNet-B0 (EFB0) [42], Inception ResNetv2 (IRv2) [39], Inception-v3 (Incv3) [40], and Inception-v4 (Incv4) [39]; and 2) Transformer-based models: ViT [10], LeViT [13], ConViT [6], Twins [5], and Pooling-based Vision Transformer (PiT) [17].
+All the models are pre-trained on the ImageNet dataset [7] .
+Additionally, we include CLIP [35], trained on 400 million text-image pairs, to evaluate the transferability across different modalities.
+When using Transformer-based models, the input image is resized to $\times224$ to meet the model input requirement.
+Source : Incv3
+Target model
+Attack RN18 RN50 VGG16 Incv3 EFB0 DN121 MBv2 IRv2 Incv4 Xcep Avg.
+DI 2.2 3.9 3.4 99.1 3.6 5.0 1.2 7.7 8.9 7.0 14.2
+RDI 5.8 5.5 3.9 99.0 8.0 8.5 3.8 18.6 18.8 11.1 18.3
+SI 6.7 6.7 4.3 98.8 9.7 9.7 4.4 23.3 22.1 13.6 19.9
+MUP 13.9 13.6 9.6 98.4 17.7 22.4 8.1 42.2 42.2 26.4 29.5
+BSR 15.8 13.9 11.9 98.7 20.5 24.3 9.6 45.7 45.5 30.3 31.6
+Admix 18.5 16.7 13.8 98.1 23.8 27.5 15.9 46.3 47.1 38.9 34.7
+SIA 17.4 21.5 16.7 98.8 27.2 32.2 13.1 56.1 59.2 42.9 38.5
+ODI 14.4 22.3 22.0 99.4 26.0 39.5 13.9 51.8 60.7 44.7 39.5
+CFM 37.4 37.9 27.3 97.9 46.1 53.0 27.8 76.9 76.1 68.2 54.9
+RaPA 51.3±0.6 53.5±1.0 41.9±0.6 97.4±0.0 60.8±0.3 68.4±1.4 48.5±0.0 86.7±0.3 87.5±0.4 84.0±0.0 68.0
+Source : ViT
+Target model
+Attack RN18 RN50 VGG16 Incv3 EFB0 DN121 MBv2 IRv2 Incv4 Xcep Avg.
+DI 0.5 1.0 0.8 2.1 2.1 1.1 1.3 1.8 1.6 1.4 1.4
+RDI 1.7 2.4 1.3 4.1 6.3 4.1 2.7 5.8 4.9 4.6 3.8
+SI 2.9 3.9 1.1 8.0 9.2 5.9 2.7 7.9 6.2 6.8 5.5
+BSR 5.0 8.2 3.9 11.2 15.0 12.8 5.3 15.5 13.4 11.2 10.2
+SE-ViT 8.9 12.0 6.9 21 25.3 20.5 8.3 23.8 22.5 20.3 17.0
+Admix 16.4 20.4 13.2 31.8 38.3 31.5 17.8 35.3 31.9 29.8 26.6
+SIA 3.6 4.8 3.0 8.1 12.2 8.4 3.5 8.9 10.1 8.3 7.1
+ODI 12.4 20.0 10.6 28.3 28.9 30.9 10.4 35.5 34.4 27.9 23.9
+CFM 26.1 33.4 18.0 45.2 56.8 47.3 23.2 54.5 49.9 46.2 40.1
+RaPA 37.2±0.7 42.9±0.7 28.6±2.0 57.5±1.4 68.2±0.4 56.8±0.2 36.4±0.7 62.9±0.4 62.6±0.1 58.3±1.0 51.2
+Table 4: ASRs (%) against ten target models on the ImageNet-Compatible dataset.
+All the attack methods are combined with MI-TI.
+The best results are shown in bold and the second best results are underlined.
+Baseline Attack Methods
+We compare RaPA with various existing transfer-based methods, including DI [52], RDI [57], SI [28], Admix [48],SIA [49], BSR [46], ODI [1], and CFM [2].
+We also include two existing self-ensemble methods, namely, MUP (whose implementation only handles CNN layers) [54] and SE-ViT (which is specifically designed for vision Transformers) [33].
+These methods are primarily used in combination with TI-FGSM [9] and MI-FGSM [8] during the optimization process.
+It is worth noting that some of these methods, namely, SI, BSR, Admix, CFM, MUP and SE-ViT, are implemented together with RDI, which have been reported to obtain higher transfer ASRs [2].
+As previously mentioned, the attack methods are configured with an identical number of inferences per iteration, denoted as $S$.
+Specifically, we pick $S$ scaled copies in SI and in the inner loop of Admix, and $S$ transformed images for BSR .
+For other baselines, we perform $S$ forward-backward passes and use the average gradient on $x_{\mathrm{adv}}$ to update the adversarial example in each iteration.
+In the main experiment, we will set $S=5$; comparison of other choices is studied in Section 4.3.
+(a) $S=1$
+(b) $S=5$
+(c) $S=10$
+Figure 2: Average ASRs along optimization iterations. Here $S$ denotes the number of inferences per iteration.
+RaPA Setting
+While the DropConnect probabilities can be chosen differently for the weight and bias parameters in a linear layer (or the transformation parameters in the normalization layer) and also across different layers, we choose the same probability for all selected parameters, that is, $p_{w}=p_{b}=p$, which greatly simplifies the implementation.
+Through our ablation study, we find that RaPA performs well across a range of probabilities.
+For our experiments, we will select the following DropConnect probabilities: $0.05$ for ResNet-50, $0.02$ for Inception-v3, $0.04$ for DenseNet-121, $0.01$ for Vision Transformer, and $0.03$ for CLIP.
+By default, RaPA applies DropConnect to all linear and normalization layers in the surrogate model.
+4.2 Main Result
+We first study the performance of RaPA on the ImageNet-Compatible dataset.
+We employ ResNet-50, Inception-v3, DenseNet-121, and ViT as surrogate models, and evaluate the obtained adversarial examples on 16 target models.
+Table 3 reports the experimental results when adversarial examples are generated using CNN-based models and transferred to Transformer-based neural networks.
+This task is considered more challenging in the context of transfer-based attacks [31], as the ASRs in this case are relatively low.
+Our method significantly improves the attack performance over existing methods: it increases the average ASR from 33.3% to 45.0% with ResNet-50 as surrogate model, and from 22.8% to 40.3% with DenseNet-121.
+Table 4 reports the ASRs of various attack methods on ten CNN-based target models.
+RaPA achieves the best average ASR.
+Particularly, with Inception-v3 as surrogate model, the ASRs are increased by 14.6% and 20.7% for the challenging target models VGG16 and MBv2, respectively.
+When transferring from Transformer-based model ViT to CNN-based models, RaPA again attains the best average ASR 51.2%.
+We also report the results of self-ensemble methods MUP [54] and SE-ViT [33].
+RaPA clearly outperforms these two methods by a large margin.
+Additional experimental results can be found in Section D.3. We also visualize the heatmaps of some adversarial examples in Appendix C for qualitative comparison.
+4.3 Ablation Study
+In this section, we conduct an ablation study to investigate the impacts of 1) different types of layers where DropConnect is applied, 2) different DropConnect probabilities and 3) more iterations and inferences.
+Different Types of Layers
+We use ResNet-50 and ViT as surrogate models to analyze the impacts of layer types.
+For ResNet-50, we apply RaPA to different combinations of Batch Normalization (BN) layer, Fully Connected (FC) layer, and Convolutional (Conv) layer.
+Similarly, we consider Layer Normalization (LN) and FC layers (including linear transformation layer in the attention layer) for ViT.
+Table 5 presents the experimental results.
+Notably, simply applying RaPA to all layers achieves equal or higher ASRs, compared with other baselines.
+The combination of BN (or LN) and FC layers performs the best, which validates our implementation in Section 3.
+For ResNet-50, applying DropConnect to Conv layers performs worse than BN layers.
+We conjecture that Conv layers have sparser weights and may be less affected by over-reliance issue.
+Applying DropConnect only to FC layers yields particularly low ASRs, as ResNet contains only a single FC layer.
+Source Model BN FC Conv ASRs (%)
+RN50 72.1
+41.1
+65.1
+72.4
+69.1
+64.7
+69.2
+Source Model LN FC - ASRs (%)
+ViT 58.6
+63.9
+65.2
+Table 5: Applying DropConnect to different types of layers. The reported ASRs(%) are averages over 16 target models.
+DropConnect Probability
+We investigate the impact of varying DropConnect probabilities $p$ using ResNet-50 as a surrogate model.
+RaPA is run with $p$ ranging from $0.01$ to $0.09$, and we report the average ASRs over 16 models in Sec.
+D.2.
+The mean ASR is 66.3% with a standard deviation of 5.9%, peaking at 72.4% when $p=0.05$.
+Notably, with $p\in[0.03,0.07]$, RaPA consistently outperforms baselines by over 2%, underscoring the stability of the proposed method across different choices.
+More Iterations and Inferences
+We study performance under different total iterations and numbers of inferences per iteration, denoted as $T$ and $S$, respectively.
+We use ResNet-50 as the surrogate model and evaluate how well adversarial examples transfer to the 16 target models.
+Figure 3: Average ASRs with different iterations ($T$) and different numbers of inferences per iteration ($S$).
+The results are reported in Figure 2.
+Although existing methods may also benefit from additional optimization iterations, RaPA and Admix have the best gains when $T$ increases, while RaPA achieves a much higher ASR than Admix.
+With $S$ increasing, RaPA can outperform CFM even at an early stage of the optimization process.
+We also depict Figure 3 to ease the comparison of the gains of different methods when both $T$ and $S$ increase.
+As we observe, RaPA benefits the most from an additional compute budget.
+4.4 Attack Performance Against Defenses
+We evaluate RaPA against several defenses: adversarially trained ResNet-50 (advRN) [36], Ensemble-Adversarial-Inception-ResNet-v2 (ensIR) [20], High-level representation Guided Denoiser (HGD) [27], Bit Depth Reduction (Bit) [15], JPEG compression [15], R&P [51], and Diffpure [34].
+We utilize ResNet-50 as the surrogate model.
+For Bit, JPEG, and R&P, the target model is ResNet-18 and for Diffpure, the target model is ResNet-50.
+As shown in Table 6, RaPA outperforms all other baselines.
+Notably, against the strong defenses ensIR and HGD, RaPA exceeds the second-best ASRs by 29.4% and 10.5%, respectively.
+RN50 Defense methods
+Method advRN ensIR JPEG Bit R&P HGD Diffpure
+DI 10.6 0.0 26.7 50.5 41.1 0.1 0.0
+RDI 39.6 0.8 58.8 75.8 71.3 0.8 0.0
+SI 61.8 9.4 75.1 80.7 78.8 2.1 0.4
+Admix 68.9 5.7 76.9 81.4 78.6 2.4 0.7
+BSR 60.4 3.0 76.8 85.2 83.4 2.6 0.1
+ODI 58.6 5.1 71.7 73.8 76.1 0.3 0.3
+CFM 84.1 13.8 88.2 91.5 89.3 15.2 0.5
+RaPA 88.2 43.2 91.2 92.2 92.7 25.7 4.0
+Table 6: ASRs (%) against six defense methods.
+Attack ViT LeViT ConViT Twins PiT CLIP Avg.
+RaPA 33.8 75.4 27.6 59.5 57.3 15.6 45.0
+DSM 8.1 49.5 7.8 31.6 23.3 3.1 20.6
+DSM-RaPA 50.8 84.2 42.2 74.6 72.6 25.1 58.3
+SASD-WS 30.1 74.1 25.6 52.1 52.9 18.1 42.2
+SASD-RaPA 42.9 79.5 35.3 63.0 63.1 25.9 51.6
+Table 7: ASRs (%) of adversarial examples generated by ResNet50 against Transformer-based targets, comparing RaPA with training-based methods (DSM and SASD-WS) and their combinations.
+Note that when RaPA is combined with SASD-WS, we did not apply Weight Scaling.
+4.5 Compatibility with Training-enhanced Frameworks
+We further compare RaPA with two representative training-required approaches, DSM[53] and SASD-WS[50], both of which involve additional optimization to enhance surrogate models for better adversarial transferability.
+Specifically, DSM trains a surrogate model with dark knowledge extracted from a teacher model and enriched by mixing augmentation, which effectively boosts the model’s transferability through knowledge-driven soft supervision.
+Meanwhile, SASD-WS improves transferability via sharpness-aware self-distillation and weight scaling, refining the loss landscape and model generalization.
+As shown in Table 7, even under a fully training-free setting, RaPA already surpasses these training-dependent methods, exhibiting superior transferability and generalization.
+Furthermore, when integrated with such training-based schemes, RaPA continues to deliver consistent gains.
+For instance, combining with DSM increases the average ASR from 20.6% to 58.3%, highlighting its excellent compatibility and scalability with existing training-enhanced frameworks.
+5 Concluding Remarks
+In this paper, we reveal the over-reliance issue in existing transfer-based attacks, where adversarial examples depend excessively on a small subset of model parameters.
+To alleviate this, we propose RaPA, which randomly prunes surrogate parameters during optimization.
+We show that the expected effect of random pruning equals adding an importance-equalization regularizer, thereby reducing parameter over-reliance and improving transferability.
+Extensive experiments on CNN and Transformer architectures confirm the effectiveness and stability of RaPA.
+Supplementary Material
+Appendix A More Related Work
 One of the most fundamental attack methods is Fast Gradient Sign Method (FGSM) [12], which uses the direction of gradient to craft adversarial examples.
@@
 Admix [48] mixes the input image with random samples from other classes.
-Clean Feature Mixup (CFM) [2] extends Admix to high-level feature space and introduces two types of competing noises to guide adversarial perturbations, achieving state-of-the-art performance in transfer-based targeted attacks.
+Clean Feature Mixup (CFM) [2] extends Admix to high-level feature space and introduces two types of competing noises to guide adversarial perturbations.
+Building upon it, Feature Tuning Mixup (FTM) [26] further introduces learnable and attack-specific feature perturbations that combine random and optimized noises in the feature space, achieving SoTA performance in transfer-based targeted attacks.
 The second class of methods stabilize the gradient updates to enhance transferability during adversarial example generation.
@@
 However, the number of surrogate models in practice is usually small and picking proper surrogate models for the same task is also not easy.
+Another line of work focuses on improving the surrogate model itself through additional training to enhance the adversarial transferability.
+Dark Surrogate Model (DSM) [53] trains a surrogate model using dark knowledge distilled from a teacher model and further augments the training data with mixing augmentations such as CutMix and Mixup, thereby enriching soft supervision and improving transferability.
+SASD-WS [50] introduces Sharpness-Aware Self-Distillation (SASD) to flatten the loss landscape and combines it with Weight Scaling (WS) to approximate model ensembling.
+While these approaches require retraining surrogate models, our RaPA works in a training-free manner, yet achieves comparable or even superior transferability, and can be seamlessly integrated with such training-based frameworks.
 Closely related to the present work is self-ensemble [19, 54, 25, 33], which creates multiple models from only one surrogate model.
-Transfer-based self-ensemble (T-sea) [18] targets the task of object detection, and the self-ensemble method in [28] specifically considers vision Transformer as surrogate model and is denoted as SE-ViT in this paper.
+Transfer-based self-ensemble (T-sea) [19] locally apply various transformations onto the input image to improve such diversity while preserving the structure of imagetargets the task of object detection, and the self-ensemble method in [33] specifically considers vision Transformer as surrogate model and is denoted as SE-ViT in this paper.
 Ghost Network [25] perturbs surrogate model to create a set of new models and then samples one model from the set at each iteration.
 Masking Unimportant Parameters (MUP) [54] drops out unimportant parameters according to a predefined Taylor expansion-based metric.
-3 Preliminary
-Consider a classification task where the model is defined as a function $f:\mathbb{R}^{n}\to\mathcal{Y}$ that maps an input $x\in\mathbb{R}^{n}$ to a label in the set $\mathcal{Y}$ consisting of all the labels.
-Given a clean image $x$ with its true label $y\in\mathcal{Y}$, untargeted attacks aim to find an adversarial example $x_{\mathrm{adv}}\in\mathbb{R}^{n}$ that is similar to $x$ but misleads the model to produce an incorrect prediction, i.e.
-$f(x_{\mathrm{adv}})\neq y$.
-Here ‘similarity’ is usually measured by an $\ell_{p}$-norm, e.g., $\|x_{\mathrm{adv}}-x\|_{p}\leq\epsilon$ where $\epsilon>0$ is a predefined perturbation budget.
-For targeted attacks, the goal is to modify the model prediction to a particular target label $y_{\mathrm{tar}}$, that is, $f(x_{\mathrm{adv}})=y_{\mathrm{tar}}\neq y$.
-In this work, we will focus on targeted attacks.
-Under the white-box setting where the model is fully accessible, adversarial example can be obtained through solving the following problem:
-$\operatorname*{arg\,max}_{x_{\mathrm{adv}}}~{}\mathcal{L}\left(f(x_{\mathrm{adv}})\right),~{}~{}\text{s.t.}~{}~{}\|x_{\mathrm{adv}}-x\|_{p}\leq\epsilon,$ (1)
-where $\mathcal{L}(\cdot)$ represents a loss function like cross-entropy loss.
-A strong and popular attack method to the above optimization problem is the Iterative Fast Gradient Sign Method (I-FGSM) [20] that leverages the sign of the gradient information.
-In particular, at each iteration $t$, the adversarial example is updated by adding a small perturbation:
-$x_{\mathrm{adv}}^{t}=x_{\mathrm{adv}}^{t-1}+\alpha\cdot\mathrm{sign}\left(\nabla_{x_{\mathrm{adv}}^{t-1}}\mathcal{L}\left(f(x_{\mathrm{adv}}^{t-1})\right)\right),$ (2)
-where $x_{\mathrm{adv}}^{0}=x$ and $\alpha>0$ is a step size.
-To make the generated adversarial example satisfy the perturbation budget constraint, a straightforward way is to project $x_{\mathrm{adv}}^{t}$ into the $\epsilon$-ball of $x$.
-In the black-box settings, the gradient information is not available and we only have limited access to the target model.
-A large class of attack methods in this setting is transfer-based, which assumes that the adversarial examples exhibit transferability across different neural models; that is, adversarial examples generated on one model to mislead not only that model but also other models [34].
-Thus, one can use the above updates (2) on an available white-box model, usually called surrogate model, to generate adversarial examples that are applied to the target model.
-To boost the performance of transfer-based methods, an effective approach is to perform some transformation $\mathcal{T}$ (like resizing [44, 48] and scaling [24]) to the input and then use $\nabla_{x_{\mathrm{adv}}^{t-1}}\mathcal{L}\left(f(\mathcal{T}(x_{\mathrm{adv}}^{t-1}))\right)$ as the gradient information.
-However, the inherent differences w.r.t.
-model architecture, parameter setting, and training procedure between the surrogate and target models still limit the effectiveness of transfer-based methods on certain models and datasets.
-In the next section, we show that there is a key aspect that renders the current over-fitting issue of transfer-based methods.
-4 Methodology
-In this section, we first conduct a pilot study to show a key aspect of the overfitting issue in existing transfer-based methods and then propose a random masking based approach.
-Comparison with related methods is also discussed.
-4.1 Motivation
-We hypothesize that the pixels or features in the adversarial example, obtained through solving Problem (1), collaborate in a highly dependent manner to maximize the success of an adversarial attack on the surrogate model, which we refer to as perturbation co-adaptation.
-Even with state-of-the-art transfer-based attack methods, this issue is still a key aspect that can make the perturbation fail to mislead other models.
-We now conduct a pilot study to verify our hypothesis.
-Since adversarial example is a composition of the given clean image and some perturbation, Problem (1) is equivalent to finding a perturbation that maximizes the loss function.
-In this sense, we perform random masking with different Dropout ratios to the obtained perturbation $\delta:=x_{\mathrm{adv}}-x$, to disrupt potential co-adaptations among pixels.
-Note that the masked version of perturbation always satisfies the $\ell_{p}$-norm constraint if the adversarial example already meets this constraint.
-We consider a targeted attack setting on an ImageNet-compatible dataset (see more details in Sec.
-5.1), where adversarial examples are crafted using ResNet-50.
-The used attack methods include DI, ODI and CFM [44, 1, 2].
-The results are shown in Fig.
-2 (additional results can be found in Sec.
-B.1).
-We can observe an obvious decrease in ASRs as the masking ratio exceeds 20%, in the DI, ODI and CFM methods, indicating that the pixels in the obtained perturbation are likely to be heavily dependent to attain a successful targeted attack.
-Similar to the co-adaption issue in training neural networks [33], we believe that perturbation co-adaptation would lead to overly complex patterns that mislead surrogate model but fail on other models.
-As such, how to further alleviate this co-adaptation would be a key to improving the transferability of adversarial examples over existing transfer-based attack methods.
-4.2 Mitigating Perturbation Co-adaptation by DropConnect
-As per the pilot study, a direct approach to improving transferability would be to apply random masking to pixels in perturbation during the iterative updates.
-As shown in Sec.
-B.2, by encouraging perturbations to explore more diverse feature disturbances, this way indeed yields higher ASRs than the baseline DI and RDI methods in the targeted attack setting.
-However, it does not consistently achieve the best attack performance; a potential reason is that pixel-level masking may be less effective of capturing semantic information for image classification.
-Based on these initial results, it is tempting to extend this direct idea to craft more transferable adversarial examples.
-[…]
-(b) ResNet-50 $\to$ MobileNet-v2
-Figure 2: ASRs(%) with increasing mask ratio.
-The rest of this work is devoted to further mitigating the perturbation co-adaptation, by extending random masking to higher-level representations.
-The overall approach is to create multiple diverse variants of the surrogate model during optimization.
-We first consider the following questions:
-  • Q1: how to achieve more flexible and diverse feature disturbances for perturbation than vanilla Dropout?
-  • Q2: which intermediate representations or layers to perform the random masking?
-For Q1, we observe that in typical machine learning problems where the goal is to optimize model parameters w.r.t.
-given training data, standard Dropout [33] is applied to the units or certain features.
-Finding adversarial example, as formulated in Problem (1), can be treated as optimizing perturbed image $x_{\mathrm{adv}}$ with fixed model parameters as ‘training data’ or ‘input features’.
-As analogy to Dropout, we can randomly drop out the model parameters connecting the units.
-This is indeed the DropConnect technique [39], a generalized version of Dropout by providing a more fine-grained way of regularizing the network during training.
-As such, we will focus on DropConnect in this paper.
-For Q2, we notice that DropConnect is mainly effective in terms of linear layers [39].
-We thus apply DropConnect to the weight and bias parameters of linear layers as well as the transformation parameters of normalization layers in the surrogate model.
-Both types of layers are widely used in mainstream architectures including Transformer [38].
-Empirically, our ablation study in Sec.
-5 validates the effectiveness of this choice, compared with convolutional layers.
-We now present our attack method, Mitigating perturbation Co-adaptation by DropConnect (MCD), as summarized in Algorithm 1.
-Take linear layer for example.
-We perform independent random masking onto the weight and bias (when present) parameters using Bernoulli sampling.
-Specifically, for surrogate model $f$, let $W\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ denote the weight matrix and $b\in\mathbb{R}^{d_{\text{out}}}$ the bias vector associated with a selected linear layer.
-Here $d_{\text{in}}$ and $d_{\text{out}}$ represent the input and output dimensions, respectively.
-The corresponding masks for the weight matrix and bias are
-$\mathcal{M}_{w}\sim\text{Bernoulli}(p_{w}),\quad\mathcal{M}_{b}\sim\text{Bernoulli}(p_{b}),$ (3)
-where $\mathcal{M}_{w}\in\{0,1\}^{d_{\text{in}}\times d_{\text{out}}}$ and $\mathcal{M}_{b}\in\{0,1\}^{d_{\text{out}}}$ are the random masks, and $p_{w},p_{b}\in[0,1]$ are DropConnect probabilities.
-Then the masked parameters are computed as
-$W_{\mathcal{M}}=\mathcal{M}_{w}\odot W,\quad b_{\mathcal{M}}=\mathcal{M}_{b}\odot b,$ (4)
-where $\odot$ denotes the element-wise multiplication.
-For normalization layer, the same operation is applied similarly to the transformation parameters.
-The random masks are sampled for each selected layer.
-Algorithm 1 Mitigating perturbation Co-adaptation by DropConnect (MCD)
-1:Classifier $f$; clean image $x$; loss function $\mathcal{L}(\cdot)$; max iterations $T$; $\ell_{p}$ bound $\epsilon$; number of inferences per iteration $S$; DropConnect probabilities $p_{w}$, $p_{b}$; linear and normalization layers $\mathbb{L}$; input transformation $\mathcal{T}$.
-2:Adversarial example $x_{\mathrm{adv}}$
-3:$x_{\mathrm{adv}}^{0}\leftarrow x$
-4:for $t=1\to T$ do
-5:     for $s=1\to S$ do
-6:         Obtain modified model $f_{\mathcal{M}}$ by applying MCD to each layer in $\mathbb{L}$ according to Eqs. (3) and (4).
-7:         $g_{s}^{t}\leftarrow\nabla_{x_{\mathrm{adv}}^{t-1}}\mathcal{L}\left(f_{\mathcal{M}}(\mathcal{T}(x_{\mathrm{adv}}^{t-1}))\right)$
-8:     end for
-9:     $g^{t}\leftarrow\frac{1}{S}\sum g_{s}^{t}$
-10:     Update $x_{\mathrm{adv}}^{t}$ with gradient $g^{t}$ using iterative methods (like MI-FGSM [7]).
-11:     Project ${x_{\mathrm{adv}}^{t}}$ into the $\epsilon$-ball of $x$.
-12:end for
-13:return $x_{\mathrm{adv}}^{T}$
-Unlike standard Dropout or DropConnect in training neural networks, MCD does not need additional re-scaling.
-The random masks $\mathcal{M}_{w}$ and $\mathcal{M}_{b}$ in Eq.
-3 are re-generated for each inference, producing diverse variants of the surrogate model with different parameters if we conduct multiple inferences at an iteration.
-In addition, MCD  can be naturally integrated with existing input transformation and gradient stabilization methods for crafting adversarial examples, as shown in Lines 5 and 8 in Algorithm 1.
-4.3 Discussion
-In this section, we compare MCD with related self-ensemble methods in more details.
-While MCD is proposed from the perspective of mitigating perturbation co-adaptation, it turns out to be a self-ensemble method that constructs multiple new models at each iteration.
-Existing self-ensemble method [18] targets object detection task, which is different from ours.
-SE-ViT [28] is specifically designed for vision Transformer surrogate models; as shown in Sec.
-5, its ASR is lower than MCD even when using ViT as surrogate model.
-More closely related are Ghost Network [22] and MUP [45], but are much outperformed by MCD (c.f.
-Sec.
-B.2 and Sec.
-5.2).
-We now analyze the effectiveness of MCD from model ensemble perspective.
-It has been hypothesized that an adversarial image that remains adversarial for multiple models is more likely to transfer to other models [26].
-MCD generates independent random masks for each selected layer and also at each optimization iteration.
-In this sense, it brings in more randomness and further diversification than Ghost Network and MUP.
-Specifically, Ghost Network perturbs only skip connections for residual networks (like ResNet-50), while MUP masks unimportant parameters according to a predefined metric at each iteration.
-This observation is also in accordance with [18, 3, 25], which show that increasing the number of surrogate models generally enhances the transfer attack performance.
-On the other hand, from the perspective of ensemble techniques in machine learning, each variant model should also be informative about or useful to the targeted task of image classification [18, 3].
-As empirically shown in Sec.
-5.5, MCD achieves a good tradeoff in terms of model diversity and utility, thereby enhancing the overall attack performance.
-In addition, MCD is simple to implement and is robust across a range of parameter choices.
-In particular, our implementation of MCD is to pick the same DropConnect probability, i.e., $p_{w}=p_{e}=p$, for selected layers.
-MCD can be applied to all layers in the model with little decrease of ASR, compared to our best practice that only selects the linear and normalization layers in this paper.
-See the experiments in Sec.
-5 for more details.
-5 Experiments
-This section conducts extensive experiments to show the effectiveness of the proposed method. In particular, we consider both CNN- and Transformer-based classifiers.
-5.1 Experimental Settings
-Dataset
-We utilize the ImageNet-compatible dataset,https://github.com/cleverhans-lab/cleverhans/tree/master/cleverhans_v3.1.0/examples/nips17_adversarial_competition/dataset which served as the official dataset for the NIPS 2017 Attack Challenge.
-This dataset contains both ground-truth and targeted labels, making it well-suited for conducting experiments on targeted attacks.
-General Setting
-We adopt the $\ell_{\infty}$-norm as the constraint on perturbation, with budget $\epsilon=16/255$.
-The learning rate is chosen as $\alpha=2/255$.
-We use the Logit loss [47] as our objective function in Eq.
-1.
-By default, we set the maximum number of optimization iterations to $1,000$ and the batch size to $32$ for all baseline methods, to ensure sufficient optimization.
-Furthermore, we notice that different attack methods may take different amounts of computations per iteration.
-We hence fix the same number of inferences for each optimization iteration to maintain fair comparisons across all attack methods.
-Source : Incv3
-Target model
-Attack RN18 RN50 VGG16 Incv3 EFB0 DN121 MBv2 IRv2 Incv4 Xcep Avg.
-DI 2.2 3.9 3.4 99.1 3.6 5.0 1.2 7.7 8.9 7.0 14.2
-RDI 5.8 5.5 3.9 99.0 8.0 8.5 3.8 18.6 18.8 11.1 18.3
-SI 6.7 6.7 4.3 98.8 9.7 9.7 4.4 23.3 22.1 13.6 19.9
-MUP 13.9 13.6 9.6 98.4 17.7 22.4 8.1 42.2 42.2 26.4 29.5
-BSR 15.8 13.9 11.9 98.7 20.5 24.3 9.6 45.7 45.5 30.3 31.6
-Admix 18.5 16.7 13.8 98.1 23.8 27.5 15.9 46.3 47.1 38.9 34.7
-ODI 14.4 22.3 22.0 99.4 26.0 39.5 13.9 51.8 60.7 44.7 39.5
-CFM 37.4 37.9 27.3 97.9 46.1 53.0 27.8 76.9 76.1 68.2 54.9
-MCD 51.9 52.5 42.1 97.5 60.6 68.6 48.7 87.1 88.1 84.1 68.1
-Source : ViT
-Target model
-Attack RN18 RN50 VGG16 Incv3 EFB0 DN121 MBv2 IRv2 Incv4 Xcep Avg.
-DI 0.5 1.0 0.8 2.1 2.1 1.1 1.3 1.8 1.6 1.4 1.4
-RDI 1.7 2.4 1.3 4.1 6.3 4.1 2.7 5.8 4.9 4.6 3.8
-SI 2.9 3.9 1.1 8.0 9.2 5.9 2.7 7.9 6.2 6.8 5.5
-BSR 5.0 8.2 3.9 11.2 15.0 12.8 5.3 15.5 13.4 11.2 10.2
-SE-ViT 8.9 12.0 6.9 21 25.3 20.5 8.3 23.8 22.5 20.3 17.0
-Admix 16.4 20.4 13.2 31.8 38.3 31.5 17.8 35.3 31.9 29.8 26.6
-ODI 12.4 20.0 10.6 28.3 28.9 30.9 10.4 35.5 34.4 27.9 23.9
-CFM 26.1 33.4 18.0 45.2 56.8 47.3 23.2 54.5 49.9 46.2 40.1
-MCD 36.2 43.9 28.2 58.0 69.0 56.7 35.8 63.2 62.9 57.8 51.2
-Table 1: ASRs (%) against ten target models on the ImageNet-Compatible dataset.
-All the attack methods are combined with MI-TI.
-The best results are shown in bold and the second best results are underlined.
-Surrogate and Target Models
-Our experiments choose various models commonly used in the literature [2].
-These include 1) CNN-based models: VGG-16 [32], ResNet-18 (RN18) [15], ResNet-50 (RN50) [15], DenseNet-121 (DN121) [17], Xception (Xcep) [4], MobileNet-v2 (MBv2) [31], EfficientNet-B0 (EFB0) [37], Inception ResNetv2 (IRv2) [36], Inception-v3 (Incv3) [35], and Inception-v4 (Incv4) [36]; and 2) Transformer-based models: ViT [9], LeViT [12], ConViT [10], Twins [5], and Pooling-based Vision Transformer (PiT) [16].
-All the models are pre-trained on the ImageNet dataset [6] .
-Additionally, we include CLIP [29], trained on 400 million text-image pairs, to evaluate the transferability across different modalities.
-When using Transformer-based models, the input image is resized to 224$\times$224 to meet the model input requirement.
-Baseline Attack Methods
-We compare MCD with various existing transfer-based methods, including DI [44], RDI [48], SI [24], Admix [42], BSR [40], ODI [1], and CFM [2].
-We also include two existing self-ensemble methods, namely, MUP (whose implementation only handles CNN layers) [45] and SE-ViT (which is specifically designed for vision Transformers) [28].
-These methods are primarily used in combination with TI-FGSM [8] and MI-FGSM [7] during the optimization process.
-It is worth noting that some of these methods, namely, SI, BSR, Admix, CFM, MUP and SE-ViT, are implemented together with RDI, which have been reported to obtain higher transfer ASRs [2].
-As previously mentioned, the attack methods are configured with an identical number of inferences per iteration, denoted as $S$.
-Specifically, we pick $S$ scaled copies in SI and in the inner loop of Admix, and $S$ transformed images for BSR .
-For other baselines, we perform $S$ forward-backward passes and use the average gradient on $x_{\mathrm{adv}}$ to update the adversarial example in each iteration.
-In the main experiment, we will set $S=5$; comparison of other choices is studied in Sec.
-5.3.
-MCD Setting
-While the DropConnect probabilities can be chosen differently for the weight and bias parameters in a linear layer (or the transformation parameters in the normalization layer) and also across different layers, we simply pick the same probability for all selected parameters, that is, $p_{w}=p_{b}=p$, which greatly simplifies the implementation.
-Through our ablation study, we find that MCD performs well across a range of probabilities.
-For our experiments, we will select the following DropConnect probabilities: $0.05$ for ResNet-50, $0.02$ for Inception-v3, $0.04$ for DenseNet-121, $0.01$ for Vision Transformer, and $0.03$ for CLIP.
-By default, MCD applies DropConnect to all linear and normalization layers in the surrogate model.
-Source: RN50 Source: DN121
-Attack ViT LeViT ConViT Twins PiT CLIP Avg. ViT LeViT ConViT Twins PiT CLIP Avg.
-DI 0.4 6.7 0.6 3.8 1.8 0.5 2.3 0.3 4.0 1.2 2.0 2.1 0.3 1.7
-RDI 2.8 24.0 4.4 12.6 10.1 1.2 9.2 1.0 12.0 2.1 6.9 8.2 1.4 5.3
-SI 8.0 42.9 7.7 25.1 23.3 3.7 18.4 3.9 21.5 4.4 10.2 14.3 2.1 9.4
-BSR 6.8 42.4 7.7 25.3 21.9 2.3 17.7 2.7 21.6 2.7 10.6 11.9 1.2 8.5
-Admix 7.2 43.4 5.9 22.7 19.4 4.2 17.1 4.2 33.8 4.0 18.0 19.1 3.9 13.8
-ODI 15.5 49.3 11.8 31.6 35.1 5.6 24.8 8.5 36.3 8.4 19.8 26.7 4.9 17.4
-MUP 7.0 48.3 8.2 30.7 26.5 3.6 20.7 5.2 38.9 5.0 23.2 20.1 4.1 16.1
-CFM 17.3 65.8 14.6 47.5 39.9 7.9 32.2 10.4 49.0 8.3 30.1 30.4 5.5 22.3
-MCD 34.5 76.7 27.1 60.1 57.1 15.7 45.2 28.0 68.5 23.2 53.8 53.6 14.1 40.2
-Table 2: ASRs (%) against five transformer-based target models on the ImageNet-Compatible dataset.
-We also include the result on CLIP.
-All the attack methods are combined with MI-TI.
-The best results are shown in bold and the second best results are underlined.
-(a) $S=1$
-(b) $S=5$
-(c) $S=10$
-Figure 3: Average ASRs along optimization iterations. Here $S$ denotes the number of inferences per iteration.
-5.2 Main Result
-We first study the performance of MCD on the ImageNet-Compatible dataset.
-We employ ResNet-50, Inception-v3, DenseNet-121, and ViT as surrogate models, and evaluate the obtained adversarial examples on 16 target models.
-Tab.
-1 reports the ASRs of various attack methods on ten CNN-based target models.
-MCD achieves the best average ASR.
-Particularly, with Inception-v3 as surrogate model, the ASRs are increased by 14.8% and 20.9% for the challenging target models VGG16 and MBv2, respectively.
-When transferring from Transformer-based model ViT to CNN-based models, MCD again attains the best average ASR of 51.2%.
-We also report the results of self-ensemble methods MUP [45] and SE-ViT [28], and MCD clearly outperforms these two methods by a large margin.
-In Tab.
-2, we present experimental results of targeted attacks, where adversarial examples are generated using CNN-based models and are transferred to Transformer-based neural networks.
-This task is considered more challenging in the context of transfer-based attacks [27], as the ASRs in this case are still relatively low.
-Again, our method significantly improves the attack performance over existing methods: it increases the average ASR from 32.2% to 45.2% with ResNet-50 as surrogate model, and from 22.3% to 40.2% when the surrogate model is DenseNet-121.
-Due to space limit, additional experimental results with different surrogate models can be found in Sec.
-B.4.
-We also visualize the heatmaps of some generated adversarial examples in Appendix A for qualitative comparison.
-5.3 More Iterations and Inferences
-We now study the performance given different total iterations and different numbers of inferences per iteration, which are denoted as $T$ and $S$, respectively.
-We use ResNet-50 as the surrogate model and evaluate how well the adversarial examples are transferred to the 16 target models.
-Figure 4: Average ASRs with different iterations (denoted by $T$) and different numbers of inferences per iteration (denoted by $S$).
-The results are reported in Fig.
-3.
-Although existing methods may also benefit from additional optimization iterations, MCD and Admix have the best gains when $T$ increases, while MCD achieves a much higher ASR than Admix.
-With $S$ increasing, MCD can outperform CFM even at an early stage of the optimization process.
-We also depict Fig.
-4 to ease the comparison of the gains of different methods when both $T$ and $S$ increase.
-As we observe, MCD benefits the most from an additional compute budget.
-5.4 Ablation Study
-In this section, we conduct an ablation study to investigate the impacts of 1) different types of layers where DropConnect is applied, and 2) different DropConnect probabilities.
-Source Model BN FC Conv ASRs (%)
-RN50 72.1
-41.1
-65.1
-72.4
-69.1
-64.7
-69.2
-Source Model LN FC - ASRs (%)
-ViT 58.6
-63.9
-65.2
-Table 3: Applying DropConnect to different types of layers. The reported ASRs(%) are averages over 16 different target models.
-Different Types of Layers
-We use ResNet-50 and ViT as surrogate models to analyze the impacts of layer types.
-For ResNet-50, we apply MCD to different combinations of Batch Normalization (BN) layer, Fully Connected (FC) layer, and Convolutional (Conv) layer.
-Similarly, we consider Layer Normalization (LN) and FC layers (including linear transformation layer in the attention layer) for ViT.
-Tab.
-3 presents the experimental results.
-Notably, simply applying MCD to all layers achieves equal or higher ASRs, compared with other baselines.
-The combination of BN (or LN) and FC layers performs the best, which validates our implementation in Sec.
-4.
-For ResNet-50, applying DropConnect to Conv layers performs worse than BN layers.
-We conjecture that Conv layers generally have sparser weights and may be less affected by perturbation co-adaptation.
-DropConnect Probability
-We further investigate the impact of varying DropConnect probabilities, denoted as $p$, using ResNet-50 as a surrogate model.
-We run MCD with different probabilities ranging from $0.01$ to $0.09$ and report the average ASRs over 16 different models.
-Along different $p$, the mean of average ASRs is 66.3% with a standard deviation of 5.9%.
-The highest average ASR is 72.4%, achieved at $p=0.05$.
-Notably, with $p\in[0.03,0.07]$, MCD consistently outperforms other baselines by more than 2%, underscoring the stability of the proposed method across different choices.
-Due to space limit, the detailed results are given in Fig.
-9 in the supplementary material.
-5.5 Model Diversity and Utility
-In this section, we show the diversity and utility of constructed models.
-Here diversity is measured by first computing the mean KL-divergence between the output distributions of each pair of model variants for an input and then averaging these values over the ImageNet-compatible dataset.
-Similarly, the utility is quantified by the mean top-1 accuracy over the model variants and the dataset.
-We use the same surrogate model ResNet-50 as in the ablation study.
-The results are presented in Fig.
-5.
-As DropConnect probability $p$ increases, the diversity among different variants improves while the utility decreases.
-As shown in Fig.
-9 in the supplementary material, the average ASR using the same surrogate model increases with $p\in[0.01,0.05]$ and then decreases with $p\in[0.05,0.09]$, while the highest ASR is obtained at $p=0.05$.
-This verifies our discussion in Sec.
-4.3 that the attack performance is affected by both diversity and utility.
-We conclude that the better performance of MCD is due to the improved diversification across the variants while keeping each variant sufficiently useful.
-5.6 Attack Performance Against Defense Methods
-Finally, we evaluate MCD against several defense mechanisms, including adversarial training methods adversarially trained ResNet-50 (advRN) [30] and Ensemble-Adversarial-Inception-ResNet-v2 (ensIR) [19], and also High-level representation Guided Denoiser (HGD) [23], Bit Depth Reduction (Bit) [14], JPEG compression [14], and R&P [43].
-We utilize ResNet-50 as the surrogate model.
-For Bit, JPEG, and R&P, the target model is ResNet-18.
-As shown in Tab.
-4, MCD  outperforms all other baselines.
-Notably, in the challenging scenario of ensIR and HGD defenses, our method surpasses the second best ASRs by 29.4% and 10.5%, respectively.
-(a) Diversity
-(b) Utility
-Figure 5: ASRs(%) with increasing mask ratio.
-RN50 Defense methods
-Method advRN ensIR JPEG Bit R&P HGD
-DI 10.6 0.0 26.7 50.5 41.1 0.1
-RDI 39.6 0.8 58.8 75.8 71.3 0.8
-SI 61.8 9.4 75.1 80.7 78.8 2.1
-Admix 68.9 5.7 76.9 81.4 78.6 2.4
-BSR 60.4 3.0 76.8 85.2 83.4 2.6
-ODI 58.6 5.1 71.7 73.8 76.1 0.3
-CFM 84.1 13.8 88.2 91.5 89.3 15.2
-MCD 88.2 43.2 91.2 92.2 92.7 25.7
-Table 4: ASRs (%) against six defense methods.
-6 Concluding Remarks
-In this paper, we first show the perturbation co-adaptation issue in existing transfer-based attack methods, and then propose MCD that creates diverse variants of surrogate model at each iteration to improve transferability.
-MCD brings in more diversification across the variants while maintaining the utility of each variant.
-Extensive experiments demonstrate the superior performance of MCD.
-Supplementary Material
-Appendix A Visualization of Adversarial Examples
+Appendix B Estimation and Approximation
+B.1 Parameter Importance Estimation and Pruning by Importance
+Computing the second-order derivative defined in Eq. (3) is generally costly. A more compact approximation is obtained via first-order expansion [32], which reduces to
+$\mathcal{I}(\theta_{i})=\frac{\partial^{2}\mathcal{L}f(x_{\mathrm{adv}})}{\partial\theta_{i}^{2}}\theta_{i}^{2}\approx\left(\frac{\partial\mathcal{L}f(x_{\mathrm{adv}})}{\partial\theta_{i}}\theta_{i}\right)^{2}.$ (7)
+In our pilot study, we adopt the method in DepGraph [11] to perform the pruning, which explicitly models inter-layer dependencies and comprehensively groups coupled parameters for pruning.
+B.2 Derivation of the Masked Loss Approximation in Eq. (4)
+Let $\theta$} represent the entire set of model parameters.
+We define a random binary mask $\mathcal{M}\in\{0,1\}^{|\theta|}$, where each entry is independently sampled from a Bernoulli distribution: $\mathcal{M}_{i}\sim\text{Bernoulli}(1-p)$.
+A second-order Taylor expansion gives
+$\mathcal{L}\bigl(f(x_{\mathrm{adv}};\mathcal{M}\odot\theta)\bigr)\approx$ $~\mathcal{L}\bigl(f(x_{\mathrm{adv}};\theta)\bigr)+\sum_{i}g_{i}\Delta_{i}+\tfrac{1}{2}\sum_{i,j}H_{ij}\Delta_{i}\Delta_{j},$
+where $\Delta=(\mathcal{M}-\mathbf{1})\odot\theta$, $g_{i}=\partial\mathcal{L}\bigl(f(x_{\mathrm{adv}};\theta)\bigr)/\partial\theta_{i}$ and $H_{ij}=\partial^{2}\mathcal{L}\bigl(f(x_{\mathrm{adv}};\theta)\bigr)/(\partial\theta_{i}\partial\theta_{j})$ evaluated at $\theta$.
+Since $p$ is taken to be small and close to zero, we have $\mathbb{E}[\Delta_{i}]\approx 0$.
+And the mask entries are independent, so we further have $\mathbb{E}[\Delta_{i}\Delta_{j}]\approx 0$ for $i\neq j$ while $\mathbb{E}[\Delta_{i}^{2}]=p(1-p)\theta_{i}^{2}$.
+Taking expectations yields
+$\mathbb{E}_{\mathcal{M}}\!\left[\mathcal{L}\bigl(f(x_{\mathrm{adv}};\mathcal{M}\odot\theta)\bigr)\right]\approx\mathcal{L}\bigl(f(x_{\mathrm{adv}};\theta)\bigr)+\frac{p(1-p)}{2}\sum_{i}H_{ii}\,\theta_{i}^{2},$ (8)
+which is the expression in Eq. (4).
+B.3 Detailed Gini Coefficients Computation Procedure
+We compute the Gini coefficients based on the parameter importance values $\mathcal{I}(\theta_{i})$ defined in Eq.
+(3).
+For each layer $l$ in the surrogate model, we first collect all parameter importance values $\{\mathcal{I}^{l}(\theta_{i})\}_{i=1}^{n}$.
+To better highlight the disparity among parameters, we further apply an exponential scaling to these importance values before computing the Gini coefficient:
+$\tilde{\mathcal{I}}^{l}(\theta_{i})=\exp\left(\mathcal{I}^{l}(\theta_{i})\right),$ (9)
+where $\tilde{\mathcal{I}}^{l}(\theta_{i})$ denotes the scaled importance.
+This transformation magnifies the relative differences among parameters, enabling a clearer assessment of importance inequality.
+Given the scaled importance values $\{\tilde{\mathcal{I}}^{l}(\theta_{1}),\tilde{\mathcal{I}}^{l}(\theta_{2}),\dots,\tilde{\mathcal{I}}^{l}(\theta_{n})\}$ in layer $l$, we compute the Gini coefficient as:
+$\text{Gini}^{l}=\frac{2\sum_{i=1}^{n}i\tilde{v}_{(i)}}{n\sum_{i=1}^{n}\tilde{v}_{(i)}}-\frac{n+1}{n},$ (10)
+where $\tilde{v}_{(i)}$ represents the $i$-th smallest value after sorting in ascending order.
+A smaller Gini coefficient indicates a more balanced distribution of parameter importance within the layer.
+Finally, we aggregate the Gini coefficients across layers to obtain both the type-wise and overall average results:
+$\text{Gini}_{\text{avg}}=\frac{1}{|\mathbb{L}|}\sum_{l\in\mathbb{L}}\text{Gini}^{l},$ (11)
+where $\mathbb{L}$ denotes the set of all layers containing learnable parameters.
+This process allows us to quantitatively evaluate the degree of importance imbalance and verify that the proposed random parameter pruning strategy effectively reduces the model’s over-reliance on a small subset of dominant parameters.
+Appendix C Visualization of Adversarial Examples
 Figs.
-6 and 7 visualize the attention heatmap of the adversarial examples, with surrogate model ResNet-50 and target model ResNet-18.
+4 and 5 visualize the attention heatmap of the adversarial examples, with surrogate model ResNet-50 and target model ResNet-18.
 We observe that the attention heatmaps w.r.t.
-MCD on the surrogate model are more similar to the heatmaps on target model, compared with other attack methods.
+RaPA on the surrogate model are more similar to the heatmaps on target model, compared with other attack methods.
 This in part explains that the generated adversarial example of MCD is more transferable to the target model.
@@
 The intensity of red indicates the level of importance assigned to each area, influencing the classifier prediction towards the target label.
-Appendix B Additional Experimental Results
-B.1 Additional Results in Pilot Study
-We present more pilot study results discussed in Sec. 4.1.
-(a) DenseNet-121 $\to$ ResNet-18
-(b) DenseNet-121 $\to$ MobileNet-v2
-Figure 8: ASRs(%) with increasing mask ratio.
-B.2 Initial Results of Applying Dropout to Perturbation and Intermediate Features
-Tab.
-5 reports the results of applying random masking to perturbation as well as to intermediate features.
-We see that the Dropout idea generally helps with increased ASR, but does not achieve the best performance.
-Notice that CFM is also implemented in combination with RDI, which we write it explicitly as RDI+CFM in the table.
+Appendix D Additional Experimental Results
+D.1 Result of masking the most important parameters
+Method DI GN RDI MMP(1%) MMP(0.1%) MMP(0.01%) MMP(0.001%)
+ASR(%) 26.8 29.7 42.0 0.2 27.7 36.7 38.9
+Table 8: Average ASRs when masking the most important parameters (MMP) in the surrogate model at mask ratios of 1 %, 0.1 %, 0.01 %, and 0.001 %.
+Results are the average over 16 target models on the ImageNet-compatible dataset, with ResNet-50 as surrogate model.
+Detailed experimental setting can be found in Section 4.1.
+Using the approximation in Eq.
+7, we mask the most important parameters.
+The results are reported in Table 8.
+As analyzed in Section 3, masking the most important parameters degrades the model’s attacking capability.
+A high masking ratio (e.g., 1 %) causes the attack to fail, whereas a low ratio (e.g., 0.001 %) leaves transferability unaffected, so the ASRs converge to the baseline RDI.
 We also include the empirical result of Ghost Networks (GN) [25].
@@
 However, its performance is still much lower.
-Compared with RDI+DropF, we can see that the additional randomness of applying Dropout to each layer is beneficial to transfer-based targeted attack.
-Method DI GN DI+DropP DI+DropF RDI RDI+DropP RDI+DropF RDI+CFM
-ASR(%) 26.8 29.7 38.7 60.0 42.0 48.7 63.0 66.1
-Table 5: Average ASRs of applying Dropout onto perturbation and intermediate features, denoted as +DropP and +DropF, respectively.
-Results are the average over 16 target models on the ImageNet-compatible dataset, with ResNet-50 as surrogate model.
-Detailed experimental setting can be found in Section 5.1.
-B.3 Ablation Study on DropConnect Probability
-Experimental results of varying DropConnect probabilities are shown in Fig. 9, indicating a stable performance of MCD across a range of different parameter values.
-Figure 9: ASRs(%) averaged over 16 different models with varying DropConnect probabilities, using ResNet-50 as the surrogate model.
-B.4 Additional Attack Results
+D.2 Ablation Study on DropConnect Probability
+Experimental results of varying DropConnect probabilities are shown in Fig. 6, indicating a stable performance of RaPA across a range of different parameter values.
+Figure 6: ASRs(%) averaged over 16 models with varying DropConnect probabilities, using ResNet-50 as surrogate .
+D.3 Additional Attack Results
 Tab.
@@
 The setting is the same as described in Sec.
-5.1.
-Our method MCD outperforms other baselines by a large margin.
-Furthermore, in the setting where CLIP is used as a surrogate model, MCD again results in a notable improvement.
+4.1.
+Our method RaPA outperforms other baselines by a large margin.
+Furthermore, in the setting where CLIP is used as a surrogate model, RaPA again results in a notable improvement.
 Yet, the ASRs are still relatively low, indicating that there is much room for further enhancement, particularly in the case of transferring from CLIP to CNN-based models.
@@
 Admix 83.2 98.1 78.1 57.1 67.8 86.8 70.3 61.9 64.4 64.5 73.2
+SIA 84.5 98.3 79.3 39.0 67.1 88.4 70.1 46.0 53.4 50.9 67.7
 ODI 77.5 98.9 83.0 63.9 69.9 89.8 61.2 71.0 71.1 69.2 75.5
@@
 CFM 92.0 98.0 89.5 74.8 85.6 93.4 85.2 82.6 82.5 81.7 86.5
-MCD 93.0 98.2 91.0 82.6 88.5 93.6 90.9 82.1 85.5 85.0 89.0
+RaPA 93.0 98.2 91.0 82.6 88.5 93.6 90.9 82.1 85.5 85.0 89.0
 Source : DN121
@@
 Admix 72.4 74.3 65.8 46.9 55.1 97.9 51.2 55.5 56.6 53.9 63.0
+SIA 69.2 79.3 66.2 34.9 54.8 98.1 45.1 40.6 51.6 44.1 58.4
 ODI 59.6 71.5 69.6 48.1 52.2 98.9 35.5 58.1 60.0 52.7 60.6
@@
 CFM 82.4 88.9 78.6 60.9 69.9 98.5 60.5 73.0 72.7 70.3 75.6
-MCD 89.6 90.2 85.1 77.8 84.4 97.4 84.4 83.2 84.1 83.7 86.0
+RaPA 89.6 90.2 85.1 77.8 84.4 97.4 84.4 83.2 84.1 83.7 86.0
 Source : CLIP
@@
 Admix 2.7 2.2 1.6 4.1 6.0 3.5 2.8 2.9 3.8 3.6 3.3
+SIA 0.9 1.0 0.8 1.8 3.5 1.6 1.9 1.3 1.8 2.6 1.7
 ODI 7.7 7.0 5.1 11.0 13.7 13.0 6.5 11.3 11.8 11.5 9.9
 CFM 3.1 3.3 1.4 4.9 7.8 4.3 2.5 5.4 4.1 4.8 4.2
-MCD 8.8 9.5 4.9 14.6 17.5 14.9 8.5 12.5 12.8 12.4 11.6
+RaPA 8.8 9.5 4.9 14.6 17.5 14.9 8.5 12.5 12.8 12.4 11.6
 Table 9: Additional experimental results against ten target models on the ImageNet-Compatible dataset, using RN50, DN121 and CLIP as surrogate model, respectively.
@@
 SI 1.0 5.0 1.0 2.0 3.8 0.4 2.2
+SIA 1.5 16.8 1.7 5.2 6.8 0.4 5.4
 BSR 1.3 15.3 1.9 5.6 7.3 0.5 5.3
@@
 CFM 9.7 43.4 8.8 23.2 26.4 3.9 19.2
-MCD 14.7 57.6 16.7 34.5 38.0 5.3 27.8
+RaPA 14.7 57.6 16.7 34.5 38.0 5.3 27.8
 Source : ViT
@@
 SI 100.0 48.2 59.2 28.3 51.4 14.8 50.3
+SIA 100.0 43.0 77.9 44.8 56.2 12.0 55.6
 BSR 100.0 57.8 63.0 45.7 67.2 18.5 58.7
@@
 CFM 99.9 90.9 94.2 82.5 91.6 55.3 85.7
-MCD 99.6 92.4 95.7 87.4 94.4 63.5 88.8
+RaPA 99.6 92.4 95.7 87.4 94.4 63.5 88.8
 Source : CLIP
@@
 SI 3.1 4.5 2.5 1.7 6.2 99.6 19.6
+SIA 7.7 7.9 5.3 3.6 8.3 99.7 22.1
 BSR 7.0 9.2 4.9 2.9 8.5 99.5 22.0
@@
 CFM 13.3 15.4 8.6 4.7 13.5 99.8 25.9
-MCD 32.5 35.5 25.1 16.7 33.0 98.4 40.2
+RaPA 32.5 35.5 25.1 16.7 33.0 98.4 40.2
 Table 10: Additional experimental results against five transformer-based target models on the ImageNet-Compatible dataset, using RN50, DN121 and CLIP as surrogate model, respectively.
@@
 The best results are shown in bold, and the second best results are underlined.
+D.4 Model Diversity and Utility
+(a) Diversity
+(b) Utility
+Figure 7: Diversity and Utility with increasing mask ratio.
+We show the diversity and utility of constructed models.
+Here diversity is measured by first computing the mean KL-divergence between the output distributions of each pair of model variants for an input and then averaging these values over the ImageNet-compatible dataset.
+Similarly, the utility is quantified by the mean top-1 accuracy over the model variants and the dataset.
+We use the same surrogate model ResNet-50 as in the ablation study.
+The results are presented in Figure 7.
+As DropConnect probability $p$ increases, the diversity among different variants improves while the utility decreases.
+As shown in Figure 6 in the supplementary material, the average ASR using the same surrogate model increases with $p\in[0.01,0.05]$ and then decreases with $p\in[0.05,0.09]$, while the highest ASR is obtained at $p=0.05$.
+This verifies our discussion in Section 3.3 that the attack performance is affected by both diversity and utility.
+We conclude that the better performance of RaPA is due to the improved diversification across the variants while keeping each variant sufficiently useful.
+Ethical Statement
+While our work is intended for research and defense purposes, it could be misused by malicious actors to craft black-box adversarial attacks against safety-critical systems such as autonomous-driving or medical models.
+By demonstrating that these systems remain vulnerable even when attackers have no knowledge of their internal details, we aim to alert service providers and researchers to this threat and to accelerate the development of more robust deep-learning defenses.